### v1

# Operator models for meromorphic functions of bounded type

Abstract
In this article, we construct operator models for meromorphic functions of bounded type on Krein spaces. This construction is based on certain reproducing kernel Hilbert spaces which are closely related to model spaces. Specifically, we show that each function of bounded type corresponds naturally to a pair of such spaces, extending Helson’s representation theorem. This correspondence enables an explicit construction of our model, where the Krein space is a suitable sum of these identified spaces. Additionally, we establish that the representing self-adjoint relations possess a relatively simple structure, proving to be partially fundamentally reducible. Conversely, we show that realizations involving such relations correspond to functions of bounded type.
1 Introduction
Quotients of bounded analytic functions, known as meromorphic functions of bounded type, were initially studied by Nevanlinna over a decade ago [41]. Since then, they have gained significant prominence in function theory, primarily due to their crucial role in the theory of Hardy spaces [44]. Nowadays, they constitute a fundamental function class in complex analysis and have applications in diverse fields such as systems theory (see e.g. [32]) and arithmetic geometry (see e.g. [29]).
Our main result establishes that functions of bounded type can be represented by operator models on Krein spaces - indefinite inner product spaces which decompose into an orthogonal sum of a Hilbert space and an anti-Hilbert space. Specifically, for a given function $q\vcentcolon U\subset\mathbb{C}^{+}\rightarrow\mathbb{C}$ of bounded type on the upper half-plane $\mathbb{C}^{+}$, we construct self-adjoint relation $A$ on a Krein space $\mathcal{K}$ such that
$U\subset\varrho(A)\quad\text{ and}\quad q(\zeta)=\overline{q(\zeta_{0})}+(\zeta-\overline{\zeta_{0}})\cdot\left[\left(I+(\zeta-\zeta_{0})(A-\zeta)^{-1}\right)v,v\right]_{\mathcal{K}},\quad\forall\zeta\in U$ (1)
$\overline{\mathrm{span}}\{\left(I+(\zeta-\zeta_{0})(A-\zeta)^{-1}\right)v\vcentcolon\zeta\in\varrho(A)\}=\mathcal{K},$ (2)
where $\zeta_{0}\in U$ is some fixed base point and $v\in\mathcal{K}$ a suitable generating vector. Such an operator model establishes a strong connection between the properties of the relation $A$ and the function $q$, for which the minimality condition (2) is essential.
Operator models have been an active field of research for at least half a decade [34]. The simplest example is given by the classical integral representation for Herglotz-Nevanlinna functions, which are analytic functions that map $\mathbb{C}^{+}$ into $\overline{\mathbb{C}^{+}}$. Any such function $q$ can be expressed as
$q(\zeta)=a+b\cdot\zeta+\int_{\mathbb{R}}\left(\frac{1}{t-\zeta}-\frac{t}{1+t^{2}}\right)\mathrm{d}\nu(t)\quad\forall\zeta\in\mathbb{\mathbb{C}}^{+},$ (3)
where $a\in\mathbb{R}$, $b\in\mathbb{R}_{\geq 0}$ and $\nu$ is a positive measure satisfying a certain regularity condition. This representation is, if suitably interpreted, a minimal realization as defined above. Indeed, if we assume $b=0$ for simplicity and choose $\mathcal{K}=L^{2}(\nu)$, $A$ to be the multiplication operator by the independent variable and $v$ appropriately, then (1) simplifies to (3). In this case, the spectrum of $A$ is given by
$\sigma(A)=\overline{\left\{x\in\mathbb{R}\vcentcolon 0<\liminf_{\epsilon\downarrow 0}\mathrm{Im}(q(x+i\epsilon))\right\}},$
which illustrates the close connection between $A$ and $q$. Finally, it turns out that all minimal realizations on Hilbert spaces are, up to an isomorphism, of this form and therefore correspond to Herglotz-Nevanlinna functions.
Consequently, in order to construct operator models for more general classes of functions, it is necessary to consider Krein spaces instead of Hilbert spaces. This more general setup allowed for the construction of minimal realizations for a wide range of functions including generalized Nevanlinna functions [35], [15], [20], definitizable functions [33], real quasi-Herglotz functions [39] and atomic density functions [22]. These functions have two things in common. First, they are of bounded type, and second, they are all closely related to Herglotz-Nevanlinna functions. For example, generalized Nevanlinna functions are products of the form $q\cdot h$, where $q$ is a Herglotz-Nevanlinna function and $h$ is a rational function satisfying a certain symmetry condition. Since the definition of “bounded type” is not related to Herglotz-Nevanlinna functions in any intrinsic way, we see that the result presented in this article is significantly more general than the preceding ones.
Finally, we mention the related result obtained in [38]. There it was shown that every function $q$ of bounded type has a non-minimal realization. This means that condition (1) is satisfied but not the minimality condition (2). However, the latter is essential to connect the function theoretic properties of $q$ with the operator theoretic properties of $A$.
Next we want to say a few words about the methods used in this article. Our construction is closely connected to the class of $\mathcal{B}(h)$-spaces that De Branges studied in a series of papers in the 1960’s [11], [12]. Here, the symbol $h$ is a meromorphic function on $\mathbb{C}\setminus\mathbb{R}$ which is bounded by $1$ on $\mathbb{C}^{+}$ and satisfies $\overline{h(\zeta)}=\frac{1}{h(\overline{\zeta})}$, and $\mathcal{B}(h)$ is the reproducing kernel Hilbert space generated by the kernel
$s_{h}(z,w)=\frac{1-h(z)\overline{h(w)}}{-i\cdot(z-\overline{w})}.$
An important feature of such a space is that the difference-quotient operator
$D_{w}\vcentcolon\mathcal{B}(h)\rightarrow\mathcal{B}(h),\quad g\mapsto\frac{g(\cdot)-g(w)}{\cdot-w}$
is bounded. Moreover, we point out that $\mathcal{B}(h)$ is (essentially) a model space if $h_{|\mathbb{C}^{+}}$ is an inner function.
Now let $f$ be a given function of bounded type, which we extend as usual by $f(\overline{\zeta})=\overline{f(\zeta)}$ to the lower half plane $\mathbb{C}^{-}$. In a first step, we show that $f$ admits a representation of the form
$f=i\cdot\frac{h_{2}-h_{1}}{h_{2}+h_{1}}\quad\text{ with }\quad\mathcal{B}(h_{1})\cap\mathcal{B}(h_{2})=\{0\},$ (4)
where $h_{1}$ and $h_{2}$ are as above. In a second step, we prove that
$\mathcal{L}(f)\vcentcolon=\frac{1}{h_{1}+h_{2}}\cdot(\mathcal{B}(h_{1})\oplus(-\mathcal{B}(h_{2}))$
is a reproducing kernel Krein space with kernel
$\quad N_{f}(\zeta,w)=\frac{f(\zeta)-\overline{f(w)}}{\zeta-\overline{w}}.$
In a third step, we show that the difference-quotient operator $D_{w}$ acts boundedly on this space. Remarkably, this follows almost directly from Leibniz’s formula. Indeed, given $g\in\mathcal{B}(h_{1})$ we calculate
$D_{w}\left(\frac{1}{h_{1}+h_{2}}\cdot g\right)=\frac{1}{h_{1}+h_{2}}\cdot D_{w}(g)+g(w)\cdot D_{w}\left(\frac{1}{h_{1}+h_{2}}\right).$
Given that point evaluations and $D_{w}$ are bounded on $\mathcal{B}(h_{1})$, it is sufficient to show that $D_{w}\left(\frac{1}{h_{1}+h_{2}}\right)\in\mathcal{L}(f)$. This computation is straightforward, and with this result, we can now establish our operator model. Specifically, we define $(A-w)^{-1}=D_{w}$ and $v=N_{f}(\cdot,\zeta_{0})$.
There are a few points to note here. First, we emphasize that only elementary facts about Krein spaces are used in our argument. Instead, we construct the space $\mathcal{L}(f)$ and the relation $A$ in very explicit ways using well-studied function spaces and Leibniz’s formula. In contrast, all of the previous realization theorems rely heavily on deep results from operator theory on Krein spaces.
Second, we draw attention to the fact that the representation provided in equation (4) extends Helson’s representation for functions with real boundary values [26, Theorem 3.1.]. When $f$ is such a function, $(h_{1})_{|\mathbb{C}^{+}}$ and $(h_{2})_{|\mathbb{C}^{+}}$ become inner functions, and the associated spaces $\mathcal{B}(h_{1})$ and $\mathcal{B}(h_{2})$ are (essentially) model spaces. In order to establish this representation, we show that two $\mathcal{B}(h)$-spaces intersect trivially if and only if the symbols are relatively prime. It’s important to note that $B(h)$-spaces are, in general, not isomorphic to de Branges-Rovnak spaces, where an analogous statement does not hold. In any case, these results in their full generality appear to be new, and we believe that they hold independent interest.
Lastly, our construction of $D_{w}$ reveals an interesting interplay between point evaluations, difference-quotients, and conjugations. We use this interplay to analyse the structure of $A$. Specifically, it is shown that there exists a fundamental decomposition $\mathcal{L}(f)=\mathcal{K}_{+}[+]\mathcal{K}_{-}$ such that the relations
$S_{+}\vcentcolon=A\cap(\mathcal{K}_{+}\times\mathcal{K}_{+})\quad\text{ and }\quad S_{-}\vcentcolon=A\cap(\mathcal{K}_{-}\times\mathcal{K}_{-})$
are closed simple symmetric operators with deficiency index $(1,1)$. Relations satisfying such a decomposition property have been extensively studied in the past [16], [17], [10]. Conversely, we show that a realization involving a relation with such a decomposition property corresponds to a function of bounded type.
This article is structured as follows. Since we use quite advanced results from various fields of complex analysis and operator theory, we decided to include an extensive preliminary section, that is Section 2. While the content of this section isn’t original in a strict sense, it remains integral to comprehending the article, as we present some well-known concepts in a new light. In Section 3, we construct minimal realizations based on the representation given by (4), with its proof to be presented in Section 4. Moving forward, in Section 5 we study the constructed relations and in Section 6 we discuss some illustrative examples.
Finally, we mention that the ideas developed in [2] had a great influence on this article. There the author studied a closely related question in the context of inverse scattering and gave some partial answers. The methods developed in this article allow us to solve that problem in full generality as well (see Section 3.1 for details).
2 Preliminaries
2.1 Krein spaces and operator models
This section aims to formally introduce operator models. We begin by discussing fundamental concepts from the theory of Krein spaces, starting with their definition [31, Section 1.4]:
Definition 2.1.
Let $\mathcal{K}$ be a complex vector space and $[\cdot,\cdot]_{\mathcal{K}}$ a sesquilinear form on $\mathcal{K}$. Then $\mathcal{K}$ is a Krein space if $\mathcal{K}$ is decomposable into a direct orthogonal sum
$\mathcal{K}=\mathcal{K}_{+}[+]\mathcal{K}_{-},$
such that $(\mathcal{K}_{+},[\cdot,\cdot]_{\mathcal{K}})$ and $-\mathcal{K}_{-}\vcentcolon=(\mathcal{K}_{-},-[\cdot,\cdot]_{\mathcal{K}})$ are Hilbert spaces.
We equip a Krein space $\mathcal{K}$ with the Hilbert space topology of $\mathcal{K}_{+}\oplus(-\mathcal{K}_{-})$, which is independent of the choice of the decomposition (see e.g. [31, Section 1.4]). The space of continuous linear operators on $\mathcal{K}$ is then denoted by $\mathscr{L}(\mathcal{K})$.
Realizations are closely related to self-adjoint relations on Krein spaces. Here, a relation is a linear subspace $A\subset\mathcal{K}\times\mathcal{K}$, its adjoint $A^{+}$ is defined as
$A^{+}=\{(u,v)\in\mathcal{K}\times\mathcal{K}\vcentcolon\forall(x,y)\in A\vcentcolon[u,y]_{\mathcal{K}}=[v,x]_{\mathcal{K}}\},$
and $A$ is self-adjoint if $A=A^{+}$. We refer to [19] for a comprehensive introduction to the theory of self-adjoint relations in the indefinite setting. Finally, realizations are defined as follows [35]:
Definition 2.2.
Let $q\vcentcolon U\subset\mathbb{C}\rightarrow\mathbb{C}$ be an analytic function. Then $q$ admits a realization on a Krein space $\mathcal{K}$ if there exists
  • a self-adjoint relation $A$ on $\mathcal{K}$
  • a base point $\zeta_{0}\in\varrho(A)\cap U$ and a generating vector $v\in\mathcal{K}$
such that
$q(\zeta)=\overline{q(\zeta_{0})}+(\zeta-\overline{\zeta_{0}})\left[\left(I+(\zeta-\zeta_{0})(A-\zeta)^{-1}\right)v,v\right]_{\mathcal{K}}\quad\forall\zeta\in\varrho(A)\cap U.$ (5)
In this case, the triple $(\mathcal{K},A,v)$ is called a realization of $q$. The realization is called minimal, if
$\overline{\mathrm{span}}\{(I+(\zeta-\zeta_{0})(A-\zeta)^{-1})v\vcentcolon\zeta\in\varrho(A)\}=\mathcal{K}.$
A minimal realization is also called operator model.
Since $A$ is self-adjoint, the resolvent set $\varrho(A)$ is symmetric with respect to $\mathbb{R}$, and $q$ satisfies $q(\overline{\zeta})=\overline{q(\zeta)}$ whenever both $\zeta$ and $\overline{\zeta}$ are in $\varrho(A)\cap U$. Thus, if $q$ is defined on a subset $U\subset\mathbb{C}^{+}$, then a realization prescribes an extension to a subset of the lower half plane $\mathbb{C}^{-}$.
Every realization induces a so called defect function, which is defined as follows (see e.g. [37]):
Definition 2.3.
Let $q\vcentcolon U\subset\mathbb{C}\rightarrow\mathbb{C}$ be an analytic function and $(\mathcal{K},A,v)$ a realization. The defect function $\phi$ is defined as
$\phi\vcentcolon\varrho(A)\rightarrow\mathcal{K},\quad\phi(\zeta)=\left(I+(\zeta-\zeta_{0})(A-\zeta)^{-1}\right)v.$
Using the resolvent identity, it can be shown that for all $\zeta,w\in\varrho(A)$ it holds that
$\begin{split}&\phi(\zeta)=(I+(\zeta-w)(A-\zeta)^{-1})\phi(w)\\ \text{ and}\quad&[\phi(\zeta),\phi(w)]_{\mathcal{K}}=\frac{q(\zeta)-\overline{q(w)}}{\zeta-\overline{w}},\ \zeta\neq\overline{w},\quad[\phi(w),\phi(\overline{w})]_{\mathcal{K}}=q^{\prime}(w),\end{split}$ (6)
where we extend $q$ to the whole of $\varrho(A)$ by the right hand side of (5) if necessary. Note that (6) means that $(\mathcal{K},A,\phi(w))$ is a realization of $q$ with base point $w$. We occasionally write $(\mathcal{K},A,\phi)$ instead of $(\mathcal{K},A,v)$ if we want to suppress the dependence on the base point $\zeta_{0}$.
Finally, two realizations $(\mathcal{K}_{i},A_{i},v_{i}),\ i\in\{1,2\}$ of the same function $q$ are isomorphic if there exists a bounded unitary operator $F\vcentcolon\mathcal{K}_{1}\rightarrow\mathcal{K}_{2}$ such that $FA_{1}F^{-1}=A_{2}$ and $F(\phi_{1}(\zeta))=\phi_{2}(\zeta)$ for all $\zeta\in\varrho(A_{1})=\varrho(A_{2})$.
2.2 Functions of bounded type
This section focuses on meromorphic functions of bounded type and their fundamental properties. In order to discuss them in a meaningful way, we begin by reviewing the Hardy spaces $\mathcal{H}^{2}(\mathbb{C}^{+})$ and $\mathcal{H}^{\infty}(\mathbb{C}^{+})$ along with inner and outer functions. A comprehensive introduction to their theory is given in [44, Section 5], where all of the upcoming definitions and results can be found.
Throughout this article, we denote by $\lambda$ the Lebesgue measure on $\mathbb{R}$ and by $\mathcal{O}(U)$ the space of analytic functions defined on an open set $U\subset\mathbb{C}$. The Hardy spaces $\mathcal{H}^{2}(\mathbb{C}^{+})$ and $\mathcal{H}^{\infty}(\mathbb{C}^{+})$ are then defined as
$\mathcal{H}^{2}(\mathbb{C}^{+})\vcentcolon$ : =$=\left\{f\in\mathcal{O}(\mathbb{C}^{+})\vcentcolon\|f\|_{\mathcal{H}^{2}(\mathbb{C}^{+})}\vcentcolon=\sup_{0<y<\infty}\left(\int_{\mathbb{R}}|f(x+iy)|^{2}\mathrm{d}\lambda(x)\right)^{\frac{1}{2}}<\infty\right\}$
$\mathcal{H}^{\infty}(\mathbb{C}^{+})\vcentcolon$ : =$=\left\{f\in\mathcal{O}(\mathbb{C}^{+})\vcentcolon\|f\|_{\mathcal{H}^{\infty}(\mathbb{C}^{+})}\vcentcolon=\sup_{\zeta\in\mathbb{C}^{+}}|f(\zeta)|<\infty\right\}.$
Both $\mathcal{H}^{2}(\mathbb{C}^{+})$ and $\mathcal{H}^{\infty}(\mathbb{C}^{+})$ are complete. Another important property is the existence of boundary values. More precisely, for any given $f\in\mathcal{H}^{2}(\mathbb{C}^{+})$ ($f\in\mathcal{H}^{\infty}(\mathbb{C}^{+})$) the limit
[…]
In fact, the same is true even for non-tangential boundary limits, but this will not be used in this article. Either way, this allows us to identify a function with its boundary function, which defines an isometric embedding $\mathcal{H}^{2}(\mathbb{C}^{+})\hookrightarrow L^{2}(\mathbb{R},\lambda)$. Consequently, $\mathcal{H}^{2}(\mathbb{C}^{+})$ is even a Hilbert space.
The elementary building blocks of functions in $\mathcal{H}^{\infty}(\mathbb{C}^{+})$ are inner and outer functions. A function $V\in\mathcal{H}^{\infty}(\mathbb{C}^{+})$ is called inner, if it satisfies $\|V\|_{\mathcal{H}^{\infty}(\mathbb{C}^{+})}\leq 1$ and $|V(\zeta)|=1$ $\lambda$-a.e. on $\mathbb{R}$. An outer function is a function of the form
$F(\zeta)\vcentcolon=\exp\left(\frac{-i}{\pi}\int_{\mathbb{R}}\left(\frac{1}{t-\zeta}-\frac{t}{1+t^{2}}\right)\log(|h(t)|)\mathrm{d}t\right),$
where $h$ is a function on $\mathbb{R}$ satisfying $h\geq 0\ \lambda$-a.e. and
$\int_{\mathbb{R}}\frac{h(t)d\lambda(t)}{1+t^{2}}<\infty.$
It can be shown that $h$ is given by $\log(|h(t)|)=|F(t)|$. Finally, for any given function $f\in\mathcal{H}^{\infty}(\mathbb{C}^{+})$, there exists an inner function $V$ and an outer function $F$ such that
$f=V\cdot F.$
This concludes our brief excursion into the theory of Hardy spaces, and we are ready to define meromorphic functions of bounded type and the Smirnov class:
Definition 2.4.
Let $f\vcentcolon U\subset\mathbb{C}^{+}\rightarrow\mathbb{C}$ be a meromorphic function on $\mathbb{C}^{+}$. Then $f$ is of bounded type if there are two functions $f_{1},f_{2}\in\mathcal{H}^{\infty}(\mathbb{C}^{+})$ such that
$f(\zeta)=\frac{f_{1}(\zeta)}{f_{2}(\zeta)}\quad\text{ for all }\zeta\in U.$
We denote this class by $\mathcal{N}(\mathbb{C}^{+})$. If $f_{2}$ can be chosen to be an outer function, then $f$ is in the Smirnov class, which we denote by $\mathcal{N}^{+}(\mathbb{C}^{+})$.
Note that $f_{1}$ and $f_{2}$ can be chosen to satisfy $\|f_{1}\|_{\mathcal{H}^{\infty}(\mathbb{C}^{+})}\leq 1$ and $\|f_{2}\|_{\mathcal{H}^{\infty}(\mathbb{C}^{+})}\leq 1$ by scaling them appropriately. Functions of bounded type inherit many properties from bounded analytic functions, for example the existence of boundary values.
Throughout this article, we will often encounter functions whose boundary functions are elements of $L^{2}(\mathbb{R})$. In that situation, the following regularity result for the Smirnov class will be useful [44, Theorem 5.23]:
Proposition 2.5.
Let $f\in\mathcal{N}^{+}(\mathbb{C}^{+})$ and assume that its boundary function is an element of $L^{2}(\mathbb{R})$. Then $f\in\mathcal{H}^{2}(\mathbb{C}^{+})$.
We emphasize that the a priori assumption $f\in\mathcal{N}^{+}(\mathbb{C}^{+})$ is essential and cannot be weakened.
Functions of bounded type and the Smirnov class on the lower-half plane $\mathbb{C}^{-}$ are defined analogously and denoted by $\mathcal{N}(\mathbb{C}^{-})$ and $\mathcal{N}^{+}(\mathbb{C}^{-})$ respectively. The spaces $\mathcal{N}(\mathbb{C}^{+})$ and $\mathcal{N}(\mathbb{C}^{-})$ are connected via the Schwartz reflection, which is the vector space isomorphism given by
$SR\vcentcolon\mathcal{N}(\mathbb{C}^{+})\rightarrow\mathcal{N}(\mathbb{C}^{-}),\quad SR(f)(z)=\overline{f(\overline{z})}.$
Note that $SR$ maps $\mathcal{H}^{2}(\mathbb{C}^{+})$ and $\mathcal{N}^{+}(\mathbb{C}^{+})$ onto $\mathcal{H}^{2}(\mathbb{C}^{-})$ and $\mathcal{N}^{+}(\mathbb{C}^{-})$ respectively, and preserves inner and outer functions.
Now recall that a realization of a function $q$ defined on $\mathbb{C}^{+}$ prescribes an extension to $\mathbb{C}^{-}$ via $q(\overline{\zeta})=\overline{q(\zeta)}$. In order to take this extension into account, we define the following function classes:
$\mathcal{N}(\mathbb{C}\setminus\mathbb{R})\vcentcolon$ : =$=\{f\vcentcolon U\subset\mathbb{C}\setminus\mathbb{R}\rightarrow\mathbb{C}\text{ meromorphic on }\mathbb{C}\setminus\mathbb{R}\vcentcolon f_{|\mathbb{C}^{+}}\in\mathcal{N}(\mathbb{C}^{+})\text{ and }f_{|\mathbb{C}^{-}}\in\mathcal{N}(\mathbb{C}^{-})\}$
$\mathcal{N}^{+}(\mathbb{C}\setminus\mathbb{R})\vcentcolon$ : =$=\{f\in\mathcal{N}(\mathbb{C}\setminus\mathbb{R})\vcentcolon f_{|\mathbb{C}^{-}}\in\mathcal{N}^{+}(\mathbb{C}^{+})\text{ and }f_{|\mathbb{C}^{-}}\in\mathcal{N}^{+}(\mathbb{C}^{-})\}$
$\mathcal{N}_{sym}\vcentcolon$ : =$=\{f\in\mathcal{N}(\mathbb{C}\setminus\mathbb{R})\vcentcolon f_{|\mathbb{C}^{-}}=SR(f_{|\mathbb{C}^{+}})\}.$
Taking the lower-half plane $\mathbb{C}^{-}$ into account may not seem important at first glance. However, this rather subtle difference turns out to be significant for reasons that will become apparent in Section 2.5.
We will often consider the Cayley transform of functions in $\mathcal{N}_{sym}$, which is defined as
$C\vcentcolon\mathcal{N}_{sym}\setminus\{f\in\mathcal{N}_{sym}\vcentcolon f_{|\mathbb{C}^{+}}\equiv i\text{ or }f_{|\mathbb{C}^{+}}\equiv-i\}\rightarrow\mathcal{N}(\mathbb{C}\setminus\mathbb{R}),\quad f\mapsto C(f)=\frac{1+if}{1-if}.$ (7)
We collect some properties in the next Proposition.
Proposition 2.6.
The Cayley transform $C$ is injective and its inverse is given by
$f=C^{-1}(C(f))=i\cdot\frac{1-C(f)}{1+C(f)}.$
Moreover, the symmetry relation $f_{|\mathbb{C}^{-}}=SR(f_{|\mathbb{C}^{+}})$ transforms to $C(f)_{|\mathbb{C}^{-}}=\frac{1}{SR(C(f)_{|\mathbb{C}^{+}})}$.
Proof.
It is straightforward to check that the inverse of $C$ is indeed given by (7). Moreover, we calculate for $\zeta\in\mathbb{C}^{-}$
$C(f)_{|\mathbb{C}^{-}}(\zeta)=\frac{1+i\cdot f_{|\mathbb{C}^{-}}(\zeta)}{1-i\cdot f_{|\mathbb{C}^{-}}(\zeta)}=\frac{1+i\cdot\overline{f_{|\mathbb{C}^{+}}(\overline{\zeta})}}{1-i\cdot\overline{f_{|\mathbb{C}^{+}}(\overline{\zeta})}}=\overline{\left(\frac{1-i\cdot f_{|\mathbb{C}^{+}}(\overline{\zeta})}{1+i\cdot f_{|\mathbb{C}^{+}}(\overline{\zeta})}\right)}=\overline{\left(\frac{1}{C(f)_{|\mathbb{C}^{+}}(\overline{\zeta})}\right)},$
where we have used $f_{|\mathbb{C}^{-}}(\zeta)=\overline{f_{|\mathbb{C}^{+}}(\overline{\zeta})}$ in the second step. This concludes the proof. ∎
Finally, let $q$ be a generalized Nevanlinna function and $(\mathcal{K},A,v)$ be a minimal realization. Then it is well known [37], that $q$ is analytically continuable through an open interval $(a,b)\subset\mathbb{R}$ if and only if $(a,b)\subset\varrho(A)$. This characterizes the resolvent set of $A$ completely in terms of the function theoretic properties of $q$. We want to prove an analogous result for functions in $\mathcal{N}_{sym}$. To this end, we give the following definition:
Definition 2.7.
Let $f\in\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$ and $U(f)$ its domain. We define $U_{ext}(f)$ as the union of $U(f)$ and all $\zeta\in\mathbb{R}$ for which there exists an open interval $\zeta\in(a,b)\subset\mathbb{R}$ such that $f$ can be analytically extended trough $(a,b)$.
2.2.1 Calculus of inner functions
Inner functions have attracted significant interest in recent decades, and their theory plays a crucial role in our construction of operator models. This is why we discuss them in detail now. As before, all results and definitions can be found in [44] and [42].
First, we recall that every inner function $V$ can be decomposed into two conceptually different inner functions, namely into a Blaschke product $B$ and a singular part $S$ [44, Theorem 5.13]. If $(\zeta_{n})$ denotes the zeros of $V$ (counting multiplicities), then the Blaschke product $B$ is defined as
$B(\zeta)$ :$\vcentcolon=\prod_{n}\exp(i\cdot\alpha_{n})\cdot\frac{\zeta-\zeta_{n}}{\zeta-\overline{\zeta_{n}}},\ \text{ where }\ \alpha_{n}\in\mathbb{R}\text{ satisfies }\exp(i\cdot\alpha_{n})\cdot\frac{i-\zeta_{n}}{i-\overline{\zeta_{n}}}\geq 0.$
Note that $B$ has the same zeros as $V$. Then $V$ can be written as
$V(\zeta)=\lambda\cdot B(\zeta)\cdot\exp\left(\frac{i}{\pi}\int_{\mathbb{R}}\left(\frac{1}{t-\zeta}-\frac{t}{1+t^{2}}\right)\mathrm{d}\nu(t)\right)\cdot\exp(i\alpha\zeta),$
where $\lambda\in\mathbb{T}$, $\nu$ is a singular Borel measure satisfying $\int_{\mathbb{R}}\frac{d\nu(t)}{1+t^{2}}<\infty$ and $\alpha\in\mathbb{R}_{\geq 0}$. An inner function given by such a exponential representation is called singular.
The class of inner functions admits an multiplicative structure, which is defined as follows [25, Definition 2.16]:
Definition 2.8.
Let $V_{1}$ and $V_{2}$ be inner functions. Then we say that
  • the function $V_{1}$ divides $V_{2}$ if there exists an inner function $V_{3}$ such that $V_{2}=V_{1}\cdot V_{3}$.
  • the functions $V_{1}$ and $V_{2}$ are relatively prime, if there exists no non-constant inner function which divides both $V_{1}$ and $V_{2}$.
It is possible to characterize when $V_{1}$ and $V_{2}$ are relatively prime in terms of the zeros of the Blaschke product and the measure of the singular part [42, Section 3.2.1]:
Proposition 2.9.
Let $V_{j}=\lambda_{j}\cdot B_{j}\cdot S_{j}\cdot\exp(i\alpha_{j}\zeta),\ j\in\{1,2\}$ be two inner functions, where $\lambda_{j}\in\mathbb{T}$, $B_{j}$ is the Blaschke product and $S_{j}$ the singular part. Then $V_{1}$ and $V_{2}$ are relatively prime if and only if
  • the Blaschke products have no common zeros,
  • either $\alpha_{1}$ or $\alpha_{2}$ is zero,
  • and the measures $\mu_{1}$ and $\mu_{2}$ representing $S_{1}$ and $S_{2}$ are mutually singular.
Finally, we discuss the boundary behaviour of inner functions. To this end, the notion of a pseudocontinuation is useful [21, Section 2]:
Definition 2.10.
Let $f\in\mathcal{N}(\mathbb{C}^{+})$ and $g\in\mathcal{N}(\mathbb{C}^{-})$. Then $f$ and $g$ are said to be pseudocontinuations of each other if their boundary functions coincide $\lambda$-a.e.
Given an inner function $V$, it is straightforward to verify that $\frac{1}{SR(V)}$ is a pseudocontinuation. The following important regularity result is an immediate consequence of Proposition 2.5.
Proposition 2.11.
Let $f\in\mathcal{H}^{2}(\mathbb{C}^{+})$ and $g\in\mathcal{N}(\mathbb{C}^{-})$. Moreover, assume that they are pseudocontinuations of each other. Then either $f\equiv g\equiv 0$ or $g\notin\mathcal{N}^{+}(\mathbb{C}^{-})$.
Proof.
The boundary function of $g$ coincides by definition with the one of $f$ and is therefore in $L^{2}(\mathbb{R})$. Consequently, if we assume that $g\in\mathcal{N}^{+}(\mathbb{C}^{-})$ then $g\in\mathcal{H}^{2}(\mathbb{C}^{-})$ by Proposition 2.5. In this case, also the Schwartz reflection $SR(g)$ is an element of $\mathcal{H}^{2}(\mathbb{C}^{+})$. Moreover, it holds that
$SR(g)(x)=\overline{g(\overline{x})}=\overline{g(x)}=\overline{f(x)}\quad\text{ for $\lambda$-a.e. $x$ in $\mathbb{R}$. }$
However, there is only one function $h\in\mathcal{H}^{2}(\mathbb{C}^{+})$ such that there exists a function $m\in\mathcal{H}^{2}(\mathbb{C}^{+})$ satisfying $h(x)=\overline{m(x)}for\ \lambda$-a.e. $x\in\mathbb{R}$, and that is the constant zero function [44, Page 112]. Therefore, we arrive at $f\equiv 0$ and $SR(g)\equiv 0$, which concludes the proof. ∎
Finally, we can also characterize when the pseudocontinuation is even an analytic continuation [42, Theorem 3.2.3]:
Proposition 2.12.
Let $V=\lambda\cdot B\cdot S\cdot\exp(i\alpha\zeta)$ be an inner function and $\mu_{s}$ the measure representing the singular part. Then $V$ can be analytically continued through an interval $(a,b)\subset\mathbb{R}$ if and only if
  • the zeros of the Blaschke product do not have an accumulation point in $(a,b)$
  • the interval $(a,b)$ does not intersect the topological support of $\mu_{s}$.
In this case, the analytic continuation coincides with the pseudocontinuation.
Recall that the topological support of a positive measure $\mu$ is defined as the set
$\{x\in\mathbb{R}\vcentcolon\mu(O)>0\text{ for every open neighbourhood $O$ of $x$}\}.$
2.3 Modeling Reproducing kernel spaces
Minimal realizations are closely connected to a specific class of reproducing kernel Krein spaces, which we will call “modeling”. In what follows, we first discuss the definition and basic properties of reproducing kernel spaces in the indefinite setting, and then explain this connection in detail. A comprehensive introduction to reproducing kernel Krein spaces is given in [30].
Definition 2.13.
Let $U\subset\mathbb{C}$ be an open set and $\mathcal{K}$ a Krein space consisting of analytic functions defined on $U$. Then $\mathcal{K}$ is a reproducing kernel Krein space if there exists a function $H\vcentcolon U\times U\rightarrow\mathbb{C}$ such that
  1. 1.
    for every $w\in U$ the function $\zeta\mapsto H(\zeta,w)$ belongs to $\mathcal{K}$.
  2. 2.
    for every $f\in\mathcal{K}$ and $w\in U$ it holds that $f(w)=[f(\cdot),H(\cdot,w)]_{\mathcal{K}}$.
Note that the kernel functions form a dense set in $\mathcal{K}$.
Similar to the Hilbert space case, various operations can be applied to reproducing kernel Krein spaces [43, Section 5]. For this article, we focus specifically on two operations: summation and conjugation.
Proposition 2.14 (Summation).
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two reproducing kernel Krein spaces on the same domain $U$, and $H_{1}$ and $H_{2}$ their respective kernels. If it holds that
$\mathcal{K}_{1}\cap\mathcal{K}_{2}=\{0\},$
then $\mathcal{K}_{1}\oplus\mathcal{K}_{2}$ is a reproducing kernel Krein space with kernel $H_{1}+H_{2}$.
Proof.
Since $\mathcal{K}_{1}\cap\mathcal{K}_{2}=\{0\}$, we can interpret elements in the direct sum $\mathcal{K}_{1}\oplus\mathcal{K}_{2}$ as analytic functions instead of pairs of analytic functions. It is then straightforward to check that
$f(w)=[f(\cdot),H_{1}(\cdot,w)+H_{2}(\cdot,w)]_{\mathcal{K}_{1}\oplus\mathcal{K}_{2}}$
for $f\in\mathcal{K}_{1}\oplus\mathcal{K}_{2}$ and $w\in U$. ∎
Proposition 2.15 (Conjugation).
Let $\mathcal{K}$ be a reproducing kernel Krein space with domain $U$ and kernel $H$. Moreover, let $g\in\mathcal{O}(U)$ be an analytic function which is not identically zero on any component of $U$. Consider the vector space $g\cdot\mathcal{K}\subset\mathcal{O}(U)$ and the operator
$M_{g}\vcentcolon\mathcal{K}\rightarrow g\cdot\mathcal{K},\quad f\mapsto g\cdot f,$
which is bijective by our assumption on $g$. We equip $g\cdot\mathcal{K}$ with the inner product that turns $M_{g}$ into an isomorphism of Krein spaces. Then $g\cdot\mathcal{K}$ is a reproducing kernel Krein space with kernel
$g(\zeta)\cdot H(\zeta,w)\cdot\overline{g(w)}.$
Proof.
The usual proof for the Pontryagin space case, which can be found in [3, Theorem 1.5.7], works exactly the same here. ∎
Finally, we discuss a common way to deal with domain issues. Let $\mathcal{K}$ be a reproducing kernel Krein space with domain $U$ and $E$ a discrete set in $U$. Then we can “forget” about the points in the set $E$ by restricting the functions in $\mathcal{K}$ to the set $U\setminus E$. This is obviously a bijective operation, which means that we end up with an isomorphic reproducing kernel Krein space $\tilde{\mathcal{K}}$, whose kernel is the restriction of the original kernel $H$ to $U\setminus E\times U\setminus E$. This is especially convenient, when we want to consider conjugations by meromorphic functions on $U$. In this case, we just exclude the poles of said function from our original domain.
2.3.1 Modeling reproducing kernel spaces
Operator models are closely related to reproducing kernel Krein spaces generated by the Nevanlinna kernel:
Definition 2.16.
Let $f\in\mathcal{N}_{sym}$ and $U_{ext}(f)$ its extended domain. The Nevanlinna kernel $N_{f}$ is then defined as
$N_{f}\vcentcolon U_{ext}(f)\times U_{ext}(f)\rightarrow\mathbb{C},\quad N_{f}(\zeta,w):=\frac{f(\zeta)-\overline{f(w)}}{\zeta-\overline{w}},\ \zeta\neq\overline{w},\quad N_{f}(w,\overline{w})=f^{\prime}(w).$
Note that $N_{f}$ has already made an appearance in Definition 2.3. The Nevanlinna kernel is closely related to the difference-quotient operator which is defined as follows:
Definition 2.17.
Let $U\subset\mathbb{C}$ be an open set and $w\in U$. The difference-quotient operator $D_{w}$ is defined as
$D_{w}\vcentcolon\mathcal{O}(U)\rightarrow\mathcal{O}(U),\quad D_{w}(f)(\zeta)=\frac{f(\zeta)-f(w)}{\zeta-w},\ \zeta\neq w,\quad D_{w}(f)(w)=f^{\prime}(w).$
It is straightforward to verify that the difference-quotient operator satisfies the following two identities:
  • (i)
    The resolvent identity: $D_{\lambda}-D_{\mu}=(\mu-\lambda)D_{\lambda}D_{\mu}$, where $\lambda,\mu\in U$.
  • (ii)
    The multiplication identity: $D_{\lambda}(f\cdot g)=f(\lambda)\cdot D_{\lambda}(g)+g\cdot D_{\lambda}(f)$, where $f,g\in\mathcal{O}(U)$.
These two concepts allow us to define the analog of a minimal realization in the realm of reproducing kernels:
Definition 2.18.
Let $f\in\mathcal{N}_{sym}$ and $U_{ext}(f)$ its extended domain. Moreover, let $\mathcal{K}$ be a reproducing kernel Krein space with kernel $N_{f}$ and domain $\Omega\subset U_{ext}(f)$, which is symmetric with respect to $\mathbb{R}$. Then $\mathcal{K}$ is called a modeling reproducing kernel space for $f$ if
$D_{w}\vcentcolon\mathcal{K}\rightarrow\mathcal{K}$
is well-defined and bounded for every $w\in\Omega$.
Every modeling reproducing kernel space gives rise to a minimal realization of a particular form, which is called canonical in the context of generalized Nevanlinna functions [15]:
Proposition 2.19.
Let $f\in\mathcal{N}_{sym}$ and $\mathcal{K}$ be a modeling reproducing kernel space for $f$ with domain $\Omega$. Then there exists a self-adjoint relation $A$ such that $(A-w)^{-1}=D_{w}$ and $\Omega\subset\varrho(A)$. Moreover, for fixed $w\in\Omega$ we can express $A$ as
$\begin{split}A&=\{(D_{w}(g),g+w\cdot D_{w}(g))\in\mathcal{K}\times\mathcal{K},\ g\in\mathcal{K}\}\\ &=\{(h,g)\in\mathcal{K}\times\mathcal{K}\vcentcolon\exists c\in\mathbb{C}\vcentcolon g(\zeta)-\zeta\cdot h(\zeta)\equiv c\}.\end{split}$ (8)
Finally, the triple $(\mathcal{K},A,\mathcal{N}_{f}(\cdot,\overline{\zeta_{0}}))$ is a minimal realization of $f$ with base point $\zeta_{0}\in\Omega$.
Proof.
Consider the operator valued function
$T\vcentcolon\Omega\rightarrow\mathscr{L}(\mathcal{K}),\quad w\mapsto D_{w}.$
The resolvent identity from Definition 2.17 can be reformulated as
$T(\lambda)-T(\mu)=(\lambda-\mu)T(\lambda)T(\mu).$
Moreover, a short calculation using the reproducing kernel property shows that
$[D_{w}(N_{f}(\cdot,z)),N_{f}(\cdot,\zeta)]_{\mathcal{K}}$ $=\left[N_{f}(\cdot,z),D_{\overline{w}}(N_{f}(\cdot,\zeta))\right]_{\mathcal{K}}\quad\forall z,\zeta\in\Omega.$
Since the kernels are dense in $\mathcal{K}$, we conclude that $T(w)=T(\overline{w})^{+}$. Consequently, it follows that there exists a self-adjoint relation $A$ such that $\Omega\subset\varrho(A)$ and
$(A-w)^{-1}=T(w)=D_{w},$
see [14, Proposition 1.1.]. Finally, it is straightforward to verify that $A$ is of the form described in equation (8).
Now pick any point $\zeta_{0}\in\Omega$ and calculate
$(I+(w-\zeta_{0})(A-w)^{-1})N_{f}(\cdot,\overline{\zeta_{0}})=N_{f}(\cdot,\overline{w}).$ (9)
Consequently, we conclude using the reproducing kernel property and $f(\overline{\zeta})=\overline{f(\zeta)}$ that
$[(I+(\zeta-\zeta_{0})(A-\zeta)^{-1})N_{f}(\cdot,\overline{\zeta_{0}}),N_{f}(\cdot,\overline{\zeta_{0}})]_{\mathcal{K}(f)}$ $=[N_{f}(\cdot,\overline{\zeta}),N_{f}(\cdot,\overline{\zeta_{0}})]_{\mathcal{K}(f)}$
$=N_{f}(\overline{\zeta_{0}},\overline{\zeta})=\frac{f(\zeta)-\overline{f(\zeta_{0})}}{\zeta-\overline{\zeta_{0}}}.$
This means that $(\mathcal{K},A,\mathcal{N}_{f}(\cdot,\overline{\zeta_{0}}))$ is indeed a realization of $f$ with base point $\zeta_{0}$. Finally, since the kernel functions are dense in $\mathcal{K}$, equation (9) ensures that this realization is also minimal. ∎
Thus any modeling reproducing kernel space is associated to a minimal realization. Conversely, any minimal realization is also associated to a modeling reproducing kernel space [22, Theorem 6.4]:
Proposition 2.20.
Let $(\mathcal{K},A,v)$ be a minimal realization of $f$ with base point $\zeta_{0}\in\varrho(A)$. Then $\mathcal{K}$ is isomorphic to a modeling reproducing kernel space $\tilde{\mathcal{K}}$ with domain $\varrho(A)$. The isomorphism is given by
$F\vcentcolon\mathcal{K}\rightarrow\tilde{\mathcal{K}},\quad x\mapsto F(x),\ F(x)(\zeta)\vcentcolon=[x,\phi(\overline{\zeta})]_{\mathcal{K}},$
where $\phi$ is the defect function of $(\mathcal{K},A,v)$.
This concludes our reformulation of (minimal) realizations in terms of reproducing kernels. Working with modeling reproducing kernel spaces is advantageous as their definition does not involve linear relations. Finally, some words on the uniqueness of operator models are in order:
Remark 2.21.
Modeling reproducing kernel spaces, and in turn minimal realizations, are in general only weakly unique. Indeed, let $\mathcal{K}_{1}(f)$ and $\mathcal{K}_{2}(f)$ be two such spaces which are defined on the same set $U$. Then the identity operator
$\mathrm{Id}\vcentcolon\mathrm{span}\{N_{f}(\cdot,w)\vcentcolon w\in U\}\subset\mathcal{K}_{1}(f)\rightarrow\mathcal{K}_{2}(f),\quad N_{f}(\cdot,w_{i})\mapsto N_{f}(\cdot,w_{i}),$
is isometric, but in general not bounded. An example illustrating this point can be found in [33, Section 2]. However, it is important to note that minimal realizations are unique up to a bounded unitary operator if $\mathcal{K}$ is a Pontryagin space [35, Remark 1.7].
2.4 Herglotz-Nevanlinna functions and realizations on Hilbert spaces
The goal of this section is to discuss Herglotz-Nevanlinna functions and the corresponding operator models on Hilbert spaces, which will play a crucial role in our construction. Their definition can be formulated in multiple equivalent ways, as detailed in [6] and the survey [40]:
Proposition 2.22.
Let $q\vcentcolon\mathbb{C}\setminus\mathbb{R}\rightarrow\mathbb{C}$ be an analytic function satisfying $q(\overline{\zeta})=\overline{q(\zeta)}$. Then the following are equivalent:
  • (A)
    The function $q$ is a Herglotz-Nevanlinna function, i.e., ${\rm Im}\,q(\zeta)\geq 0$ for ${\rm Im}\,\zeta>0$.
  • (B)
    There exist numbers $a\in\mathbb{R}$ and $b\geq 0$ and a positive Borel measure $\nu$ satisfying
    $\int_{\mathbb{R}}\frac{d\nu(t)}{1+t^{2}}<\infty$
    such that
    $q(\zeta)=a+b\cdot\zeta+\int_{\mathbb{R}}\left(\frac{1}{t-\zeta}-\frac{t}{1+t^{2}}\right)\mathrm{d}\nu(t)\quad\forall\zeta\in\mathbb{\mathbb{C}}\setminus\mathbb{R}.$ (10)
  • (C)
    The Nevanlinna kernel $N_{f}$ is positive definite.
Moreover, if $q$ is a Herglotz-Nevanlinna function, then the absolutely continuous part of $\nu$ is given by
$\mathrm{d}\nu_{abs}(t)=\frac{1}{\pi}\lim_{\epsilon\downarrow 0}\mathrm{Im}(q(t+i\epsilon))\mathrm{d}\lambda(t)=\frac{1}{2i\cdot\pi}\cdot\left(\lim_{\epsilon\downarrow 0}q(t+i\epsilon)-\lim_{\epsilon\downarrow 0}q(t-i\epsilon)\right)$
see [45, Theorem 3.27]. We denote the class of Herglotz-Nevanlinna functions by $\mathcal{N}_{0}$.
It is well known that Herglotz-Nevanlinna functions are of bounded type, i.e. that $\mathcal{N}_{0}\subset\mathcal{N}_{sym}$. The integral representation in item (B) defines a minimal realization in the following way (see e.g. [36, Section 2.4]):
Proposition 2.23.
Let $q$ be a Herglotz-Nevanlinna function, $U_{ext}(q)$ its extended domain and $(a,b,\nu)$ as in Proposition 2.22. We define a Hilbert space $\mathbb{C}_{b}$ as
$\begin{cases}\mathbb{C}_{b}=\mathbb{C},\quad[\cdot,\cdot]_{\mathbb{C}_{b}}\vcentcolon=b\cdot[\cdot,\cdot]_{\mathbb{C}}\quad&\text{ if }b\neq 0\\ \mathbb{C}_{b}=\{0\}\quad&\text{ if }b=0.\end{cases}$
Then the following holds:
  • If $b=0$, then $q$ has a minimal realization $(\mathcal{K},A,\phi)$ given by
    $\mathcal{K}=L^{2}(\mathbb{R},\nu)$
    $A=\{(f,g)\in L^{2}(\mathbb{R},\nu)\times L^{2}(\mathbb{R},\nu)\ \big{|}\ g(t)\equiv t\cdot f(t)\}$
    $\phi(\zeta)=\frac{1}{t-\zeta}\in L^{2}(\mathbb{R},\nu).$
  • If $b>0$, then $q$ has a minimal realization given by
    $\mathcal{K}=L^{2}(\mathbb{R},\nu)\oplus\mathbb{C}_{b},$
    $A=\left\{\left(\begin{pmatrix}f\\ c\end{pmatrix},\begin{pmatrix}g\\ d\end{pmatrix}\right)\in L^{2}(\mathbb{R},\nu)\oplus\mathbb{C}_{b}\times L^{2}(\mathbb{R},\nu)\oplus\mathbb{C}_{b}\ \big{|}\ g(t)\equiv t\cdot f(t),\ c=0\right\},$
    $\phi(\zeta)=\begin{pmatrix}\frac{1}{t-\zeta}\\ 1\end{pmatrix}\in L^{2}(\mathbb{R},\nu)\oplus\mathbb{C}_{b}.$
The relation $A$ is an operator if and only if $b=0$. In either case it holds that $\varrho(A)=U_{ext}(q)$.
Consequently, we can use Proposition 2.20 to define a (unique) modelling reproducing kernel Hilbert space for Herglotz-Nevanlinna functions as it was done in e.g. [11, Theorem 5]:
Proposition 2.24.
Let $q$ be a Herglotz-Nevanlinna function. Then $q$ has a modeling reproducing kernel space $\mathcal{L}(q)$, which is described by the isomorphism
$F\vcentcolon\ L^{2}(\mathbb{R},\nu)\oplus\mathbb{C}_{b}\rightarrow\mathcal{L}(q),\quad x\mapsto F(x),\quad F(x)(\zeta)=[x,\phi(\overline{\zeta})]_{L^{2}(\mathbb{R},\nu)\oplus\mathbb{C}_{b}}.$
In particular, it holds that
[…]
is the sum of a Cauchy transform and a constant, which means that $\mathcal{L}(q)\subset\mathcal{N}^{+}(\mathbb{C}\setminus\mathbb{R})$, see e.g. [9, Proposition 5.1.1.].
2.5 $\mathcal{B}(h)$-spaces
As outlined in the introduction, we base our construction on a generalization of Helson‘s representation theorem for real complex functions. While it was sufficient to work with inner functions for that specific class, we need to work with the class $\mathcal{S}_{0}$ in the general case:
Definition 2.25.
Let $h\in\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$. Then $h$ is in the class $\mathcal{S}_{0}$ if
$h_{|\mathbb{C}^{+}}\in\mathcal{H}^{\infty}(\mathbb{C}^{+})\setminus\{0\},\quad\|h_{|\mathbb{C}^{+}}\|_{\mathcal{H}^{\infty}(\mathbb{C}^{+})}\leq 1,\quad\text{ and }h_{|\mathbb{C}^{-}}=\frac{1}{SR(h_{|\mathbb{C}^{+}})}.$
Note that $h$ has a pole at $\zeta\in\mathbb{C}^{-}$ if and only if $h(\overline{\zeta})=0$.
It is clear that any given function $f\in\mathcal{H}^{\infty}(\mathbb{C}^{+})\setminus\{0\}$ satisfying $\|f\|_{\mathcal{H}^{\infty}(\mathbb{C}^{+})}\leq 1$ generates a function $\tilde{f}$ in $\mathcal{S}_{0}$ by setting
$\tilde{f}=f\ \text{ on $\mathbb{C}^{+}$ and }\ \tilde{f}_{|\mathbb{C}^{-}}=\frac{1}{SR(f_{|\mathbb{C}^{+}})}.$
In this case, we say that $\tilde{f}$ is generated by $f$. It should be noted that if $f$ is an inner function, then its extension to $\mathbb{C}^{-}$ coincides with its pseudocontinuation.
The connection between the class of Nevanlinna functions and the class $\mathcal{S}_{0}$ is given by the Cayley transform. Indeed, it is straightforward to check that
$C\vcentcolon\mathcal{N}_{0}\setminus\{q\in\mathcal{N}_{0}\vcentcolon q_{|\mathbb{C}^{+}}\equiv i\}\rightarrow\mathcal{S}_{0}\setminus\{h\equiv-1\},\quad q\mapsto C(q)=\frac{1+iq}{1-iq}$
is a bijection.
The class $\mathcal{S}_{0}$ was originally introduced by de Branges [8] and has been studied on multiple occasions afterwards [2],[4]. Naturally associated with a function in $\mathcal{S}_{0}$ is the reproducing kernel Hilbert space described in the next Proposition:
Theorem 2.26 (DeBranges).
Let $h\in\mathcal{S}_{0}$. Then there exists a unique reproducing kernel Hilbert space with domain $U_{ext}(h)$ and kernel
$s_{h}(z,w)\vcentcolon U_{ext}(h)\times U_{ext}(h)\rightarrow\mathbb{C}$
$s_{h}(z,w)=\frac{1-h(z)\overline{h(w)}}{-i\cdot(z-\overline{w})},\ z\neq\overline{w},\quad s_{h}(w,\overline{w})=\frac{1}{-i\cdot h(w)}\cdot h^{\prime}(w).$
We denote this space by $\mathcal{B}(h)$. Moreover, the difference-quotient operator $D_{w}$ is well-defined and bounded on $\mathcal{B}(h)$ for every $w\in U_{ext}(h)$ such that $h(w)\neq 0$.
A proof can be found in [4, Section 3]. Alternatively, one can proceed as follows:
Theorem 2.27.
Let $h\in\mathcal{S}_{0}\setminus\{h\equiv-1\}$ and define $q=C^{-1}(h)$. Then the reproducing kernel space given by
$\frac{1+h(\zeta)}{\sqrt{2}}\cdot\mathcal{L}(q)\subset\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$
is the space $\mathcal{B}(h)$, as described in Theorem 2.26. Furthermore, for any $w\in U_{ext}(h)$ such that $h(w)\neq 0$ it holds that
$D_{w}(h)=\frac{h(\zeta)-h(w)}{\zeta-w}\in\mathcal{B}(h).$
Proof.
We calculate
$\frac{1+h(\zeta)}{\sqrt{2}}\cdot$ $N_{q}(\zeta,w)\cdot\frac{1+\overline{h(w)}}{\sqrt{2}}=\frac{1}{2\cdot(\zeta-\overline{w})}\cdot(1+h(\zeta))\cdot(q(\zeta)-\overline{q(w)})\cdot(1+\overline{h(w)})$
$=\frac{1}{2\cdot(\zeta-\overline{w})}\cdot(1+h(\zeta))\cdot\left(i\cdot\frac{1-h(\zeta)}{1+h(\zeta)}+i\cdot\frac{1-\overline{h(w)}}{1+\overline{h(w)}}\right)\cdot(1+\overline{h(w)})$
$=\frac{1}{2\cdot-i\cdot(\zeta-\overline{w})}\cdot((1-h(\zeta))\cdot(1+\overline{h(w)})+(1+h(\zeta))\cdot(1-\overline{h(w)}))=s_{h}(\zeta,w).$
Consequently, we see that the Nevanlinna kernel $N_{q}$ and $s_{h}$ only differ by a conjugation. According to Proposition 2.15, this means that
$\mathcal{B}(h)=\frac{1+h}{\sqrt{2}}\cdot\mathcal{L}(q)\subset\mathcal{N}(\mathbb{C}\setminus\mathbb{R}).$ (11)
Next, let $w\in U_{ext}(h)$ such that $h(w)\neq 0$. Since $h$ is analytic at $\overline{w}$ and $h(w)\cdot\overline{h(\overline{w})}=1$ by construction, we conclude that
$-D_{w}(h)(\zeta)=-\frac{h(\zeta)-h(w)}{\zeta-w}=-i\cdot h(w)\cdot\frac{1-h(\zeta)\cdot\overline{h(\overline{w})}}{-i\cdot(\zeta-w)}=h(w)\cdot s_{h}(\zeta,\overline{w})\in\mathcal{B}(h).$ (12)
Finally, consider $f\in\mathcal{B}(h)$. We can express $f$ as $f=(h+1)\cdot\tilde{f}$, where $\tilde{f}\in\mathcal{L}(q)$. Using the multiplication formula for $D_{w}$, we infer that
$D_{w}(f)=(h+1)\cdot D_{w}(\tilde{f})+\tilde{f}(w)\cdot D_{w}(h+1)=(h+1)\cdot D_{w}(\tilde{f})+\tilde{f}(w)\cdot D_{w}(h).$
Here, we have used the fact that the difference quotient doesn’t “see” constants. Since $D_{w}$ is a bounded operator on $\mathcal{L}(q)$ and $D_{w}(h)\in\mathcal{B}(h)$, it follows that $D_{w}(f)\in\mathcal{B}(h)$. By the closed graph theorem $D_{w}$ is also bounded, which completes the proof. ∎
A careful reading of the proof reveals that the boundedness of $D_{w}$ is equivalent to $D_{w}(h)\in\mathcal{B}(h)$. We will discuss this relationship at a later stage. Next, we clarify the relationship between $\mathcal{B}(h)$ spaces and the more widely known de Branges-Rovnak spaces, which are defined as follows:
Definition 2.28.
Let $f\in\mathcal{H}^{\infty}(\mathbb{C}^{+})$ with $\|f\|_{\mathcal{H}^{\infty}(\mathbb{C}^{+})}\leq 1$ and define the kernel function
$k_{f}\vcentcolon\mathbb{C}^{+}\times\mathbb{C}^{+}\rightarrow\mathbb{C},\quad k_{f}(\zeta,w)=\frac{1-f(\zeta)\cdot\overline{f(w)}}{-i\cdot(\zeta-\overline{w})}.$
This kernel generates a reproducing kernel Hilbert space $\mathcal{H}(f)$, called de Branges-Rovnak space or model space if $f$ is an inner function. Finally, it holds that $\mathcal{H}(f)\subset\mathcal{H}^{2}(\mathbb{C}^{+})$ as sets.
This raises the question, when the restriction to $\mathbb{C}^{+}$ is an isomorphism from $\mathcal{B}(h)$ to $\mathcal{H}(h_{|\mathbb{C}^{+}})$, which is answered e.g. in [1, Section 6]:
Proposition 2.29.
Let $h\in\mathcal{S}_{0}$ be extreme, which means that
$\int_{\mathbb{R}}\frac{\log(1-|\lim_{\epsilon\downarrow 0}h(x+i\epsilon)|)}{1+x^{2}}\mathrm{d}x=-\infty.$
Then the restriction operator
$\mathrm{Res}\vcentcolon\mathcal{B}(h)\rightarrow\mathcal{H}(h_{|\mathbb{C}^{+}}),\quad f\mapsto f_{|\mathbb{C}^{+}}$
is an isometric isomorphism.
We see that there is a subtle difference between the two sided space $\mathcal{B}(h)$ and the one sided de Branges-Rovnak space. This difference will turn out to be crucial, which is why take the lower half plane into account and consider functions in $\mathcal{N}_{sym}$ instead of $\mathcal{N}(\mathbb{C}^{+})$.
Next, we note that the class $\mathcal{S}_{0}$ has a multiplicative structure. Specifically, if $h_{1}$ and $h_{2}$ are elements of $\mathcal{S}_{0}$, then so is $h_{1}\cdot h_{2}$. This enables us to establish a divisibility theory analogous to the one for inner functions (Section 2.2):
Definition 2.30.
Let $h_{1}$ and $h_{2}$ in $\mathcal{S}_{0}$. Then we say that
  • the function $h_{1}$ divides $h_{2}$ if there exists a function $h_{3}\in\mathcal{S}_{0}$ such that $h_{2}=h_{1}\cdot h_{3}$.
  • the functions $h_{1}$ and $h_{2}$ are relatively prime, if there exists no non-trivial function in $\mathcal{S}_{0}$ which divides both $h_{1}$ and $h_{2}$. Here a function $h\in\mathcal{S}_{0}$ is non-trivial, if $h_{|\mathbb{C}^{+}}$ is not a constant function of modulus one.
Similarly to the case of inner functions, we can characterize when two functions are relatively prime:
Proposition 2.31.
Let $h_{1}$ and $h_{2}$ be elements of $\mathcal{S}_{0}$ and let $(h_{i})_{|\mathbb{C}^{+}}=V_{i}\cdot O_{i},\ i\in\{1,2\}$ be the inner outer decomposition. Then $h_{1}$ and $h_{2}$ are relatively prime if and only if
  • the inner functions $V_{1}$ and $V_{2}$ are relatively prime
  • and the absolutely continuous measures $\log|O_{1}|\cdot d\lambda$ and $\log|O_{2}|\cdot d\lambda$ on $\mathbb{R}$ representing $O_{1}$ and $O_{2}$ are mutually singular.
Proof.
It is straightforward to verify that $h_{1}$ and $h_{2}$ are relatively prime if $h_{1}$ and $h_{2}$ satisfy (i) and (ii).
Conversely, let $h_{1}$ and $h_{2}$ be elements of $\mathcal{S}_{0}$, and assume that $h_{1}$ and $h_{2}$ are relatively prime. If $V_{1}$ and $V_{2}$ are not relatively prime, then we can find an inner function $V_{3}$ that divides both $V_{1}$ and $V_{2}$. In this case, $h_{3}$ generated by $V_{3}$ will divide both $h_{1}$ and $h_{2}$, leading to a contradiction.
Now, consider the situation where $\log|O_{1}|\leq 0$ and $\log|O_{2}|\leq 0$ representing $O_{1}$ and $O_{2}$ are not mutually singular. In this case, we define the outer function
$O_{3}(\zeta)\vcentcolon=\exp\left(\frac{-i}{\pi}\int_{\mathbb{R}}\left(\frac{1}{t-\zeta}-\frac{t}{1+t^{2}}\right)\cdot\min\{\log|O_{1}|,\log|O_{2}|\}\mathrm{d}t\right),$
Since for $\lambda$-a.e. $x\in\mathbb{R}$ it holds that $\log|O_{3}(x)|=\min\{\log|O_{1}(x)|,\log|O_{2}(x)|\}\leq 0$ , we conclude that $\|O_{3}\|_{\mathcal{H}^{\infty}(\mathbb{C}^{+})}\leq 1$. The function $h_{3}\in\mathcal{S}_{0}$ generated by $O_{3}$ will then divide both $h_{1}$ and $h_{2}$, leading to a contradiction once again. This concludes the proof. ∎
Finally, we will need the following description of the boundary behaviour of functions in $\mathcal{B}(h)$.
Proposition 2.32.
Let $h\in\mathcal{S}_{0}$ and set
$P_{h}=\left\{x\in\mathbb{R}\vcentcolon\lim_{\epsilon\downarrow 0}h(x+i\epsilon)=\lim_{\epsilon\downarrow 0}h(x-i\epsilon)\right\}.$
Then for every $g\in\mathcal{B}(h)$ it holds that
$\lim_{\epsilon\downarrow 0}g(x+i\epsilon)=\lim_{\epsilon\downarrow 0}g(x-i\epsilon)\quad\text{ for $\lambda$-a.e. $x\in P_{h}$}.$
Proof.
Since the space $\mathcal{B}(-1)$ is trivial, we can assume without loss of generality that $h\not\equiv-1$. This allows us to consider the Herglotz-Nevanlinna function $q$ defined as the inverse Cayley transform of $h$, i.e.
$q=C^{-1}(h)=i\cdot\frac{1-h}{1+h}.$ (13)
Let $\nu$ be the representing measure of $q$ given by Proposition 2.22. Then we can represent an arbitrary element $g\in\mathcal{B}(h)$ as
[…]
for some $\tilde{g}\in L^{2}(\mathbb{R},\nu)$ and $c\in\mathbb{C}$. The boundary limits from above and below clearly coincide for $h+1$ and constants on $P_{h}$. Therefore, we only have to treat the Cauchy-integral
[…]
Now if $\tilde{g}$ is a non-negative function, then $m$ is a Herglotz-Nevanlinna function with representing measure $\tilde{g}d\nu$ (see [40, Theorem 2.4.2]). Then we conclude that (Proposition 2.22)
$(\tilde{g}\mathrm{d}\nu)_{abs}(t)=\frac{1}{2i\cdot\pi}\cdot\left(\lim_{\epsilon\downarrow 0}m(t+i\epsilon)-\lim_{\epsilon\downarrow 0}m(t-i\epsilon)\right).$
The same formula is also true for arbitrary functions $g\in L^{2}(\mathbb{R},d\nu)$. In that case, we decompose $g$ into its positive and negative real and imaginary parts and apply the same argument after pulling out $i$ and minus signs if necessary. Finally, we calculate the absolutely continuous part of $d\nu$ using Proposition 2.22
$(d\nu)_{abs}(t)$ $=\frac{1}{2i\cdot\pi}\cdot\left(\lim_{\epsilon\downarrow 0}q(t+i\epsilon)-\lim_{\epsilon\downarrow 0}q(t-i\epsilon)\right)$
$=\frac{1}{2i\cdot\pi}\cdot\left(\lim_{\epsilon\downarrow 0}i\cdot\frac{1-h(t+i\epsilon)}{1+h(t+i\epsilon)}-\lim_{\epsilon\downarrow 0}i\cdot\frac{1-h(t-i\epsilon)}{1+h(t-i\epsilon)}\right).$
This means that $d\nu_{abs}(t)=0$ on $P_{h}$ and consequently $(\tilde{g}\mathrm{d}\nu)_{abs}(t)=0\ \lambda\text{ a.e. on }\mathbb{R}$, which concludes the proof. ∎
2.6 Boundary triples and symmetric operators
In this last preliminary section, we discuss symmetric operators and their boundary mappings, which have been extensively studied in the past [14], [7]. These concepts will play no role in the actual construction of minimal realizations for functions of bounded type, but they will be crucial for characterizing the appearing relations afterward. As always, we start with the definition:
Definition 2.33.
Let $S$ be a closed operator on a Krein space $\mathcal{K}$. Then $S$ is called symmetric if $S\subset S^{+}$.
Boundary mappings are then defined as follows:
Definition 2.34.
Let $S$ be a symmetric operator on a Krein space $\mathcal{K}$. A triple $(\mathcal{H},\Gamma_{0},\Gamma_{1})$ is called boundary triple for $S^{+}$, if $\mathcal{H}$ is a Hilbert space and $\Gamma_{0}$ and $\Gamma_{1}$ are bounded mappings from $S^{+}$ into $\mathcal{H}$ such that
  • the mapping $\Gamma_{0}\times\Gamma_{1}$ into $\mathcal{H}\times\mathcal{H}$ is surjective
  • the abstract Lagrange identity holds:
    $[g,h^{\prime}]_{\mathcal{K}}-[h,g^{\prime}]_{\mathcal{K}}=(\Gamma_{1}((h,g)),\Gamma_{0}((h^{\prime},g^{\prime})))_{\mathcal{H}}-(\Gamma_{0}((h,g)),\Gamma_{1}((h^{\prime},g^{\prime})))_{\mathcal{H}}\quad\forall(h,g),(h^{\prime},g^{\prime})\in S^{+}.$
Naturally associated to such a boundary triple is the self-adjoint linear relation $A\vcentcolon=\mathrm{ker}(\Gamma_{0})$. In what follows, we will always assume that $\varrho(A)\neq\emptyset$ and define the defect subspaces
$\mathcal{F}_{w}\vcentcolon$ : =$=\{g\in\mathcal{K}\ |\ (g,w\cdot g)\in S^{+}\}=\mathrm{ker}(S^{+}-w)$
~$\tilde{\mathcal{F}}_{w}\vcentcolon$ : =$=\{(g,w\cdot g)\ |\ g\in\mathcal{F}_{w}\}\subset\mathcal{K}\times\mathcal{K}.$
The relation $S^{+}$ can be decomposed as $S^{+}=A\oplus\tilde{\mathcal{F}}_{z}$ for $z\in\varrho(A)$, which can be used to show that
$(\Gamma_{0})_{|\tilde{\mathcal{F}}_{z}}\vcentcolon\tilde{\mathcal{F}}_{z}\rightarrow\mathcal{H}$
is a bijective map. This allows us to define the following objects:
Definition 2.35.
Let $S$ be a symmetric operator on a Krein space $\mathcal{K}$, $(\mathcal{H},\Gamma_{0},\Gamma_{1})$ a boundary triple and assume that $\varrho(A)\neq\emptyset$, where $A=\mathrm{ker}(\Gamma_{0})$. Then we define
  • the Weyl solution $\gamma\vcentcolon\varrho(A)\rightarrow\mathscr{L}(\mathcal{H},\mathcal{K})$ as
    $\gamma(z)\vcentcolon=P_{1}\left((\Gamma_{0})_{|\tilde{\mathcal{F}}_{z}}\right)^{-1}\vcentcolon\mathcal{H}\rightarrow\mathcal{K},$
    where $P_{1}$ denotes the projection onto the first component of $\mathcal{K}\times\mathcal{K}$. This function satisfies
    $\gamma(w)=(I+(w-z)(A-w)^{-1})\gamma(z)\quad\forall w,z\in\varrho(A).$
  • the Weyl function $m\vcentcolon\varrho(A)\rightarrow\mathscr{L}(\mathcal{H})$ as
    $m(z)\vcentcolon=\Gamma_{1}\left((\Gamma_{0})_{|\tilde{\mathcal{F}}_{z}}\right)^{-1}.$
If $\mathcal{K}$ is a Hilbert space, then the dimension of $\mathcal{F}_{w}$ is constant on $\mathbb{C}^{+}$ and $\mathbb{C}^{-}$ which allows us to define the deficiency index as $(n,m)=(\mathrm{dim}(\mathcal{F}_{i}),\mathrm{dim}(\mathcal{F}_{-i}))$. Boundary triples on Hilbert spaces are closely related to realizations of Herglotz-Nevanlinna functions [10, Section 2]:
Proposition 2.36.
Let $q$ be a Herglotz-Nevanlinna function and $\mathcal{L}(q)$ its modeling reproducing kernel space. Then the multiplication operator by the independent variable, which we denote by $S_{q}$, is a closed, simple symmetric operator with deficiency index $(1,1)$. Its adjoint is given by
$S^{+}=\{(f,g)\in\mathcal{L}(q)\times\mathcal{L}(q)\ \big{|}\exists c,d\in\mathbb{C}\vcentcolon g(\zeta)-\zeta\cdot f(\zeta)=c-d\cdot q(\zeta)\}.$
Finally, a boundary triple $(\mathbb{C},\Gamma_{0},\Gamma_{1})$ is defined by $\Gamma_{0}=c$ and $\Gamma_{1}=d$ and the associated Weyl function is $q$.
Recall that a symmetric operator on a Hilbert space $\mathcal{H}$ is simple or completely non-self adjoint, if there exists no non-trivial orthogonal decomposition $\mathcal{H}=\mathcal{H}_{1}\oplus\mathcal{H}_{2}$ such that $S$ decomposes as $S=S_{1}\oplus S_{2}$ and $S_{2}$ is self adjoint in $\mathcal{H}_{2}$. This Proposition has the following inverse [10, Section 2]:
Proposition 2.37.
Let $S$ be a closed, simple symmetric operator on a Hilbert space $\mathcal{H}$ with deficiency index $(1,1)$. Then there exists a boundary triple $(\mathbb{C},\Gamma_{0},\Gamma_{1})$. Moreover, given such a boundary triple, then the Weyl function $q$ is a Herglotz-Nevanlinna function and the operator
$F\vcentcolon\mathcal{H}\rightarrow\mathcal{L}(q),\quad x\mapsto[x,\gamma(\overline{\zeta})]_{\mathcal{H}}\quad\zeta\in\mathbb{C}\setminus\mathbb{R}$
is an isometric isomorphism such that $S=F^{-1}S_{q}F$.
We point out, that for fixed $x\in\mathcal{H}$ the function $\zeta\mapsto[x,\gamma(\overline{\zeta})]$ is an element of $\mathcal{L}(q)$ and thus in particular of $\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$. This will be important later on.
3 Minimal realizations for functions of bounded type
Now we turn to the proof of our main theorem:
Theorem 3.1.
Let $f\in\mathcal{N}_{sym}$. Then there exists a modeling reproducing kernel space $\mathcal{L}(f)$ with domain $U_{ext}(f)$ for $f$. In other words, $f$ has a minimal realization $(\mathcal{L}(f),A,v)$ with $\varrho(A)=U_{ext}(f)$.
As outlined in the introduction, we base our construction on a suitable pair of $\mathcal{B}(h)$ spaces. In order to do so, we establish a connection between a function $f\in\mathcal{N}_{sym}$ and two functions $h_{1},h_{2}\in\mathcal{S}_{0}$ via the Cayley transform:
Proposition 3.2.
Let $f\in\mathcal{N}_{sym}$ and assume that $f$ is non-constant on $\mathbb{C}^{+}$. Then there exist functions $h_{1},h_{2}\in\mathcal{S}_{0}$ which are relatively prime such that
$f=i\cdot\frac{h_{2}-h_{1}}{h_{2}+h_{1}}.$
Proof.
Let $f\in\mathcal{N}_{sym}$ be such that $f_{|\mathbb{C}^{+}}$ is non-constant, and consider its Cayley transform $C(f)\in\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$. We can express $C(f)_{|\mathbb{C}_{+}}\in\mathcal{N}(\mathbb{C}^{+})$ as
$C(f)_{|\mathbb{C}_{+}}=\frac{V_{1}}{V_{2}}\cdot F=\frac{V_{1}}{V_{2}}\cdot\exp\left(\frac{-i}{\pi}\int_{-\infty}^{\infty}\frac{1}{t-\zeta}-\frac{t}{1+t^{2}}\frac{1}{1+t^{2}}\cdot m(x)\mathrm{d}\lambda\right)$
where $V_{1}$ and $V_{2}$ are inner functions which are relatively prime and $F$ is the outer factor. The absolutely continuous measure $m(x)\mathrm{d}\lambda$ can be decomposed into its positive and negative part, i.e. into $m_{+}(x)\mathrm{d}\lambda$ and $m_{-}(x)\mathrm{d}\lambda$, which are mutually singular. We define outer functions $O_{1}$ and $O_{2}$ as
$O_{1}(\zeta)$ $=\exp\left(\frac{-i}{\pi}\int_{-\infty}^{\infty}\left(\frac{1}{t-\zeta}-\frac{t}{1+t^{2}}\right)\cdot m_{-}(x)\mathrm{d}\lambda\right)$
$O_{2}(\zeta)$ $=\exp\left(\frac{-i}{\pi}\int_{-\infty}^{\infty}\left(\frac{1}{t-\zeta}-\frac{t}{1+t^{2}}\right)\cdot(-m_{+}(x))\mathrm{d}\lambda\right).$
Note that $\|V_{i}\cdot O_{i}\|_{\mathcal{H}^{\infty}(\mathbb{C}^{+})}\leq 1$, since $\log|O_{1}(x)|=m_{-}(x)\leq 0$ and $\log|O_{2}(x)|=-m_{+}(x)\leq 0$. Moreover, by construction it holds that
$C(f)_{|\mathbb{C}_{+}}=\frac{V_{1}\cdot O_{1}}{V_{2}\cdot O_{2}}.$
We define $h_{1}$ and $h_{2}$ as the elements in $\mathcal{S}_{0}$ generated by $V_{i}\cdot O_{i}$. The functions $h_{1}$ and $h_{2}$ are relatively prime by construction. And since $C(f)_{|\mathbb{C}^{-}}=\frac{1}{SR(C(f)_{|\mathbb{C}^{+}})}$ (Proposition 2.6), it follows that
$C(f)=\frac{h_{1}}{h_{2}}\quad\text{ or equivalently }\quad f=C^{-1}(C(f))=C^{-1}\left(\frac{h_{1}}{h_{2}}\right)=i\cdot\frac{h_{2}-h_{1}}{h_{2}+h_{1}},$
which concludes the proof. ∎
In summary, every function $f\in\mathcal{N}_{sym}$ can be represented in terms of two functions in $\mathcal{S}_{0}$ which are relatively prime. This allows us to generalize Helson‘s representation theorem for functions with real boundary values:
Theorem 3.3.
Let $h_{1}$ and $h_{2}$ be elements of $\mathcal{S}_{0}$. Then $h_{1}$ and $h_{2}$ are relatively prime if and only if
$\mathcal{B}(h_{1})\cap\mathcal{B}(h_{2})=\{0\}.$
In particular, every function $f\in\mathcal{N}_{sym}$ which is non-constant on $\mathbb{C}^{+}$ has a representation of the form
$f=i\cdot\frac{h_{2}-h_{1}}{h_{2}+h_{1}}\quad\text{ and }\quad\mathcal{B}(h_{1})\cap\mathcal{B}(h_{2})=\{0\}.$
De Branges studied this intersection problem in the 1960s and (essentially) showed the “only if” direction [11, Theorem 6]. The “if” direction, which is the more difficult one, will be proven in Section 4. In that section, we will also introduce the following technical lemma crucial for addressing complexities associated with the real line.
Lemma 3.4.
Let $h_{1}$ and $h_{2}$ be functions in $\mathcal{S}_{0}$ which are relatively prime. Then the function
$\frac{h_{1}}{h_{2}}$
can be analytically continued through an open interval $(a,b)\subset\mathbb{R}$ if and only if both $h_{1}$ and $h_{2}$ can be analytically continued through $(a,b)$.
For the moment, let’s assume the preceding Theorem and Lemma to be true and proceed to construct a modeling reproducing kernel space based on these results. We are now almost ready to do so. However, there is a minor issue with the representation of a function $f\in\mathcal{N}_{sym}$ as
$f=i\cdot\frac{h_{2}-h_{1}}{h_{2}+h_{1}}.$
Indeed, the function $f$ remains analytic even if $h_{1}$ or $h_{2}$ has a pole, in which case $f$ takes the values $i$ or $-i$. The poles (and zeros as well) of $h_{1}$ and $h_{2}$ will require additional consideration. To address these exceptional points, we will use the following Lemma:
Lemma 3.5.
Let $f\in\mathcal{N}_{sym}$ and $(\mathcal{K},A,v)$ be a minimal realization of $f$. Moreover, fix $\zeta_{0}\in\mathbb{C}$. If there exists an $\epsilon>0$ such that $B_{\epsilon}(\zeta_{0})\subset U_{ext}(f)$ and $B_{\epsilon}(\zeta_{0})\setminus\{\zeta_{0}\}\subset\varrho(A)$, then $\zeta_{0}\in\varrho(A)$.
This result is very well-known in the Pontryagin space case, see e.g. [18, Theorem 1.1]. In [22, Section 6] the same proof was shown to work in the Krein space setting under the additional assumption that $f$ has real boundary values. However, this assumption on $f$ was not used in the proof, which is why we take Lemma 3.5 for granted.
Proof of Theorem 3.1.
Let $f\in\mathcal{N}_{sym}$. If $f$ is constant on $\mathbb{C}^{+}$, then either $f$ or $-f$ is a Herglotz-Nevanlinna function and Theorem 3.1 is clearly true. Consequently, we can assume without loss of generality that this is not the case, which allows us to find $h_{1}$ and $h_{2}$ as described in Proposition 3.2. Consider the reproducing kernel Hilbert spaces
$\mathcal{K}_{+}\vcentcolon=\frac{\sqrt{2}}{h_{1}+h_{2}}\cdot\mathcal{B}(h_{1})\subset\mathcal{N}(\mathbb{C}\setminus\mathbb{R}),\quad\mathcal{K}_{-}\vcentcolon=\frac{\sqrt{2}}{h_{1}+h_{2}}\cdot\mathcal{B}(h_{2})\subset\mathcal{N}(\mathbb{C}\setminus\mathbb{R}),$
where the multiplication is regarded as a conjugation (Proposition 2.15). We want to take the sum of these spaces, for which we first specify a suitable common domain. It is clear that both $\mathcal{K}_{+}$ and $\mathcal{K}_{-}$ consists of functions which are at least analytic on the set
$\Omega(f)\vcentcolon=\{\zeta\in\mathbb{C}\vcentcolon h_{1}\text{ and }h_{2}\text{ are analytic at $\zeta$, }h_{1}(\zeta)+h_{2}(\zeta)\neq 0\text{ and }h_{1}(\zeta)\neq 0\neq h_{2}(\zeta)\}.$
The zeros of $h_{1}$ and $h_{2}$ were excluded as well in order to ensure that $D_{w}$ acts boundedly on $\mathcal{B}(h_{i})$ for every $w\in\Omega(f)$ (see Theorem 2.26). Note that $f$ is given by
$f=i\cdot\frac{h_{2}-h_{1}}{h_{2}+h_{1}},$
which means that $\Omega(f)\subset U_{ext}(f)$. Henceforth, we consider $\mathcal{K}_{+}$ and $\mathcal{K}_{-}$ as reproducing kernel spaces consisting of functions defined on $\Omega(f)$ by restricting the domains accordingly. We have seen in Proposition 3.2, that the intersection of $\mathcal{K}_{+}$ and $\mathcal{K}_{-}$ is trivial. Therefore, their direct sum $\mathcal{L}(f)\vcentcolon=K_{+}\oplus K_{-}$ together with the inner product
$[f_{+}+f_{-},g_{+}+g_{-}]_{\mathcal{L}(f)}\vcentcolon=[f_{+},g_{+}]_{\mathcal{K}_{+}}-[f_{-},g_{-}]_{\mathcal{K}_{-}}\quad\forall f_{+},g_{+}\in\mathcal{K}_{+},\quad f_{-},g_{-}\in\mathcal{K}_{-}.$
is a reproducing kernel Krein space. Its kernel is $N_{f}(\zeta,w)$, because
$\frac{\sqrt{2}}{h_{1}(\zeta)+h_{2}(\zeta)}$ $\cdot(s_{h_{1}}(\zeta,w)-s_{h_{2}}(\zeta,w))\cdot\frac{\sqrt{2}}{\overline{h_{1}(w)}+\overline{h_{2}(w)}}$
$=\frac{\sqrt{2}}{h_{1}(\zeta)+h_{2}(\zeta)}\cdot\left(\frac{1-h_{1}(\zeta)\cdot\overline{h_{1}(w)}}{-i\cdot(\zeta-\overline{w})}-\frac{1-h_{2}(\zeta)\cdot\overline{h_{2}(w)}}{-i\cdot(\zeta-\overline{w})}\right)\cdot\frac{\sqrt{2}}{\overline{h_{1}(w)}+\overline{h_{2}(w)}}$
$=\frac{i}{\zeta-\overline{w}}\cdot\left(\frac{2\cdot(h_{2}(\zeta)\cdot\overline{h_{2}(w)}-h_{1}(\zeta)\cdot\overline{h_{1}(w)})}{(h_{1}(\zeta)+h_{2}(\zeta))\cdot(\overline{h_{1}(w)}+\overline{h_{2}(w)})}\right)$
$=\frac{i}{\zeta-\overline{w}}\cdot\left(\frac{h_{2}(\zeta)-h_{1}(\zeta)}{h_{1}(\zeta)+h_{2}(\zeta)}+\frac{\overline{h_{2}(w)}-\overline{h_{1}(w)}}{\overline{h_{1}(w)}+\overline{h_{2}(w)}}\right)=\frac{f(\zeta)-\overline{f(w)}}{\zeta-\overline{w}}=N_{f}(\zeta,w).$
Next, we show that the difference quotient operator is well-defined and bounded on this space. To this end, fix $w\in\Omega(f)$ and pick some arbitrary element $g\in\mathcal{B}(h_{1})$. This corresponds to an element
$\frac{1}{h_{1}+h_{2}}\cdot g\in\mathcal{K}_{+}.$
We calculate using the multiplication formula for the difference quotient
$D_{w}\left(\frac{1}{h_{1}+h_{2}}\cdot g\right)=\frac{1}{h_{1}+h_{2}}\cdot D_{w}(g)+g(w)\cdot D_{w}\left(\frac{1}{h_{1}+h_{2}}\right).$
Since $D_{w}$ is a bounded operator on $\mathcal{B}(h_{1})$ we conclude that $\frac{1}{h_{1}+h_{2}}\cdot D_{w}(g)\in\mathcal{L}(f)$. Moreover, it it is easy to see that
$D_{w}\left(\frac{1}{h_{1}+h_{2}}\right)=-\frac{1}{h_{1}+h_{2}}\cdot(D_{w}(h_{1})+D_{w}(h_{2}))\cdot\frac{1}{h_{1}(w)+h_{2}(w)}$ (14)
and we conclude that $D_{w}\left(\frac{1}{h_{1}+h_{2}}\right)\in\mathcal{L}(f)$ by Theorem 2.27. The same argument shows that also $D_{w}(g)\in\mathcal{L}(f)$ for $g\in\mathcal{K}_{-}$. Consequently, the operator $D_{w}$ is everywhere defined on $\mathcal{L}(f)$ and is closed since point evaluations are continuous. This means that $D_{w}$ is even continuous by the closed graph theorem.
In summary, we have constructed a modeling reproducing kernel space on $\Omega(f)$, or equivalently a minimal realization $(\mathcal{L}(f),A,\mathcal{N}_{f}(\cdot,\overline{\zeta_{0}}))$ with $\Omega(f)\subset\varrho(A)$ and $\zeta_{0}\in\Omega(f)$. Now let $\zeta\in U_{ext}(f)$. Then the following holds:
  • If $\zeta\in\mathbb{C}\setminus\mathbb{R}$ such that $\zeta\notin\Omega(f)$, then either $h_{1}$ or $h_{2}$ has a zero or pole at $\zeta$.
  • If $\zeta\in\mathbb{R}$, then $\zeta\in\Omega(f)$. Indeed, consider an interval $\zeta\in(a,b)$ on which $f$ is analytic. Then $f$ is real valued there. Consequently, it follows that $C(f)$ given by
    $C(f)=\frac{1+i\cdot f}{1-i\cdot f}=\frac{h_{1}}{h_{2}}$
    is analytic on $(a,b)$ and does not take the values $-1$ nor 00 there. This implies that $h_{1}$ and $h_{2}$ are analytic on $(a,b)$ (Theorem 3.3) and that $h_{i}\neq 0$ and $h_{1}+h_{2}\neq 0$ on $(a,b)$.
Thus, it follows that $\Omega(f)$ falls short of $U_{ext}(f)$ only by a set of isolated points, namely the poles and zeros of $h_{1}$ and $h_{2}$. We apply Lemma 3.5 to each of these points to conclude that $U_{ext}(f)=\varrho(A)$. This means that every function in $\mathcal{L}(f)$ can be analytically extended to $U_{ext}(f)$ and that $D_{w}$ is well defined and bounded on $U_{ext}(f)$. This concludes the proof. ∎
The existence of the analytic continuation onto the exceptional points might seem mysterious at first glance since it followed from an abstract argument. However, this can also be proven directly in the following way:
Remark 3.6.
Let $q_{1}$ be the Cayley transform of $h_{1}$, which is a Herglotz-Nevanlinna function. Then it follows that
$\mathcal{K}_{+}=\frac{1}{h_{1}+h_{2}}\cdot\mathcal{B}(h_{1})=\frac{h_{1}+1}{h_{1}+h_{2}}\cdot\mathcal{L}(q_{1}).$
From this equality it can be read off that any function $f\in\mathcal{K}_{+}$ has an analytic continuation at $\zeta$ even if $\zeta$ is a pole of $h_{1}$. Indeed the function
$\frac{h_{1}+1}{h_{1}+h_{2}}$
is analytically continuable at such a pole and $\mathcal{L}(q_{1})$ consists of functions which are analytic on $\mathbb{C}\setminus\mathbb{R}$.
3.1 The key argument revisited
The argument made in the proof of Theorem 3.1 is a special case of the following general principle:
Theorem 3.7.
Let $\mathcal{K}$ be a reproducing kernel Krein space on a domain $\Omega$, and assume that $D_{w}$ is well defined and bounded for every $w\in\Omega$. Furthermore, let $f\in\mathcal{O}(\Omega)$ such that
  • (i)
    the function $D_{w}(f)$ is an element of $\mathcal{K}$
  • (ii)
    the function $\frac{1}{f}$ is meromorphic on $\Omega$.
In this case, the difference-quotient operator $D_{w}$ is well-defined and bounded on the reproducing kernel Krein space $\frac{1}{f}\cdot\mathcal{K}$ for every $w\in\Omega\setminus\{w\in\Omega\vcentcolon f(w)=0\}$. Note that we do not assume that $f\in\mathcal{K}$.
Proof.
Let $\frac{1}{f}\cdot g\in\frac{1}{f}\cdot\mathcal{K}$ and calculate
$D_{w}\left(\frac{1}{f}\cdot g\right)=\frac{1}{f}\cdot D_{w}(g)+g(w)\cdot D_{w}\left(\frac{1}{f}\right).$
Since point evaluations are continuous and $\frac{1}{f}\cdot D_{w}(g)\in\frac{1}{f}\cdot\mathcal{K}$ by assumption, we conclude that $D_{w}$ is well defined and bounded on $\frac{1}{f}\cdot\mathcal{K}$ if and only if
$D_{w}\left(\frac{1}{f}\right)\in\frac{1}{f}\cdot\mathcal{K},\text{ which is equivalent to }f\cdot D_{w}\left(\frac{1}{f}\right)=\frac{1}{f(w)}\cdot D_{w}(f)\in\mathcal{K}.$
The latter statement is true by assumption. ∎
This principle allows us to define realizations of Möbius transformations. More precisely, let $q$ be of bounded type and $\mathcal{L}(f)$ be a modeling reproducing kernel space. Then $\frac{1}{f}\cdot\mathcal{L}(f)$ is a reproducing kernel Krein space with kernel
$\frac{1}{f(\zeta)}\cdot N_{f}(\zeta,w)\cdot\frac{1}{\overline{f(w)}}=N_{\frac{1}{-f}}(\zeta,w).$
Moreover, $D_{w}$ is bounded on $\frac{1}{f}\cdot\mathcal{L}(f)$ since $D_{w}(f)=N_{f}(\cdot,\overline{w})\in\mathcal{L}(f)$, which means that this is a modeling reproducing kernel space for $\frac{1}{-f}$ This was known before [38], however this reasoning significantly simplifies the proof given there.
In the following, we describe yet another application of this principle. So far we were interested in reproducing kernel spaces associated to the Nevanlinna kernel generated by functions in $\mathcal{N}_{sym}$. In the context of inverse scattering theory [2], one often considers reproducing kernel spaces corresponding to the kernel
$s_{v}(\zeta,w)=\frac{1-v(\zeta)\overline{v(w)}}{-i\cdot(\zeta-\overline{w})}.$
In this case, the natural symmetry condition for the symbol $v\in\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$ is
$v_{|\mathbb{C}^{-}}=\frac{1}{SR(v_{|\mathbb{C}^{+}})}.$
Of course, if $v_{|\mathbb{C}^{+}}$ is analytic and bounded by one, then $v\in\mathcal{S}_{0}$ and we end up with a $\mathcal{B}(v)$ space. However, the methods developed in this article allow us to define such spaces for arbitrary symbols of bounded type, and furthermore show that they are invariant under the difference-quotient operator:
Theorem 3.8.
Let $v\in\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$ and assume that $v$ satisfies the symmetry condition
$v_{|\mathbb{C}^{-}}=\frac{1}{SR(v_{|\mathbb{C}^{+}})}.$
Then there are functions $h_{1}$, $h_{2}$ in $\mathcal{S}_{0}$ such that $v=\frac{h_{1}}{h_{2}}$ and $\mathcal{B}(h_{1})\cap\mathcal{B}(h_{2})=\{0\}$. Moreover, the space
$\mathcal{B}(v)\vcentcolon=\frac{1}{h_{2}}\cdot(\mathcal{B}(h_{1})\oplus-\mathcal{B}(h_{2}))$
is a reproducing kernel Krein space with
domain $U_{0}(v)\vcentcolon=\{\zeta\in\mathbb{C}\vcentcolon h_{1}\text{ and }h_{2}\text{ are analytic at }\zeta\text{ and }h_{1}(\zeta)\neq 0\neq h_{2}(\zeta)\}$
and kernel $s_{v}(\zeta,w)\vcentcolon U_{0}(v)\times U_{0}(v)\rightarrow\mathbb{C},\quad s_{v}(\zeta,w)=\frac{1-v(\zeta)\overline{v(w)}}{-i\cdot(\zeta-\overline{w})}.$
Finally, $D_{w}$ is well defined and bounded for every $w\in U_{0}(v)$.
Proof.
It is clear that we can find $h_{1}$ and $h_{2}$ as described above. Then the space $\mathcal{B}(h_{1})\oplus-\mathcal{B}(h_{2})$ is a reproducing kernel Krein space consisting of functions which are analytic at least on $U_{0}$. The kernel of $\mathcal{B}(v)$ is $s_{v}(\cdot,w)$, because
$s_{v}(\cdot,w)=\frac{1}{h_{2}(\cdot)}\cdot(s_{h_{1}}(\cdot,w)-s_{h_{2}}(\cdot,w))\cdot\frac{1}{\overline{h_{2}(w)}}.$
Finally, the $D_{w}$ is well defined and bounded on this space, because $D_{w}$ is well defined and bounded on $\mathcal{B}(h_{1})\oplus-\mathcal{B}(h_{2})$ and $D_{w}(h_{2})\in\mathcal{B}(h_{1})\oplus\mathcal{B}(h_{2})$, see equation (12). Note that we have used Theorem 3.7 here. ∎
If $v_{|\mathbb{C}^{+}}$ has unimodular boundary values, then the space $\mathcal{B}(v)$ was considered in [2, Section 6] for the first time. However, they did not show that the difference quotient operator is bounded.
4 The proof of Theorem 3.3 and Lemma 3.4
4.1 Intersections of $\mathcal{B}(h)$-spaces
In this section, we give the missing proofs for Theorem 3.3 and Lemma 3.4. We begin with the intersection result, which is restated for convenience here:
Proposition 4.1.
Let $h_{1}$ and $h_{2}$ be elements of $\mathcal{S}_{0}$. Then $h_{1}$ and $h_{2}$ are relatively prime if and only if
$\mathcal{B}(h_{1})\cap\mathcal{B}(h_{2})=\{0\}.$
First, we treat the case in which both $h_{0}$ and $h_{1}$ are generated by an inner function. If that is the case, then we have seen in Proposition 2.29 that $\mathcal{B}(h)$ is isomorphic to the model space associated to $h_{|\mathbb{C}^{+}}$. More precisely, the following holds:
Proposition 4.2.
Let $h,h_{2}\in\mathcal{S}_{0}$ be generated by an inner functions $V$ and $V_{2}$ respectively. Then the following holds:
  • (i)
    The restriction operator $\mathrm{R}\vcentcolon\mathcal{B}(h)\rightarrow\mathcal{H}(V)\subset\mathcal{H}^{2}(\mathbb{C}^{+})$ is an isometric isomorphism. Moreover, for every $g\in\mathcal{B}(h)$ the functions $g_{|\mathbb{C}^{+}}$ and $g_{|\mathbb{C}^{-}}$ are pseudocontinuations of each other.
  • (ii)
    Assume that $V$ and $V_{2}$ are relatively prime. Then it holds that $\mathcal{H}(V)\cap\mathcal{H}(V_{2})=\{0\}$ and consequently also
    $\mathcal{B}(h)\cap\mathcal{B}(h_{2})=\{0\}.$
  • (iii)
    Let $g\in\mathcal{B}(h)\setminus\{0\}$. Then there exists a function $F\in\mathcal{H}^{2}(\mathbb{C}^{-})$ such that
    $g_{|\mathbb{C}^{-}}=\frac{F}{SR((V)_{|\mathbb{C}^{+}})},\quad F\in\mathcal{H}^{2}(\mathbb{C}^{-}).$
    Moreover, $g_{|\mathbb{C}^{-}}$ is not an element of $\mathcal{N}^{+}(\mathbb{C}^{-})$.
Proof.
Statement (i) is just a reformulation of [25, Proposition 7.14] and statement (ii) can be found in [25, Corollary 5.9]. Finally, statement (iii) can be found in [27, Page 43]. Note that $g_{|\mathbb{C}^{-}}$ cannot be an element of $\mathcal{N}^{+}(\mathbb{C}^{-})$ because of Proposition 2.11. ∎
Now we will turn to the general case, in which the outer factors no longer have to be trivial. This is exactly the point at which it is important to take the lower half plane $\mathbb{C}^{-}$ into account. Indeed, much of what will be said below is only true for $\mathcal{B}(h)$ spaces and not for the respective De Branges-Rovnak spaces. We first prove an intersection result for purely outer functions:
Lemma 4.3.
Let $h_{1},h_{2}\in\mathcal{S}_{0}$ be generated by outer functions $O_{1}$ and $O_{2}$. If $h_{1}$ and $h_{2}$ are relatively prime, then
$\mathcal{B}(h_{1})\cap\mathcal{B}(h_{2})=\{0\}.$
Proof.
If either $h_{1}$ or $h_{2}$ is trivial, then either $\mathcal{B}(h_{1})$ or $\mathcal{B}(h_{2})$ is trivial, and in this case there is nothing to show. Therefore, we can assume without loss of generality that both $h_{1}$ and $h_{2}$ are not trivial, which means that both $O_{1}$ and $O_{2}$ are non-constant.
Let $g\in\mathcal{B}(h_{1})\cap\mathcal{B}(h_{2})$ and recall that the measures $\log(|O_{1}|))\cdot d\lambda$ and $\log(|O_{2}|)\cdot d\lambda$ on $\mathbb{R}$ are mutually singular. We prove that $g$ has to be zero in three steps:
1. The functions $g_{|\mathbb{C}^{+}}$ and $g_{|\mathbb{C}^{-}}$ are pseudocontinuations of each other.
First, we observe that for $\lambda$-a.e. $x\in\mathbb{R}$
[…]
Indeed, the boundary limits of $O_{1}$ and $O_{2}$ exist $\lambda$-almost everywhere and are smaller or equal $1$ $\lambda$-a.e. by construction. Moreover, if there existed a set $E\subset\mathbb{R}$ with positive measure on which both $|O_{1}|$ and $|O_{2}|$ are strictly smaller than $1$ simultaneously, then the densities $\log(|O_{1}|)$ and $\log(|O_{2}|)$ would not be mutually singular.
Now fix a point $x\in\mathbb{R}$ such that $\lim_{\epsilon\downarrow 0}O_{j}(x+i\epsilon)\in\mathbb{T}$ for some $j\in\{1,2\}$. Then $x$ belongs to set $P_{h_{j}}$, which was introduced in Proposition 2.32 . To demonstrate this, recall that $P_{h_{j}}$ was defined as
$P_{h_{j}}=\left\{x\in\mathbb{R}\vcentcolon\lim_{\epsilon\downarrow 0}h_{j}(x+i\epsilon)=\lim_{\epsilon\downarrow 0}h_{j}(x-i\epsilon)\right\}.$
Moreover, note that $O_{j}$ coincides with $(h_{j})_{|\mathbb{C}^{+}}$. Then we calculate
$\lim_{\epsilon\downarrow 0}h_{j}(x-i\epsilon)=\frac{1}{\lim_{\epsilon\downarrow 0}\overline{h_{j}(x+i\epsilon)}}=\lim_{\epsilon\downarrow 0}h_{j}(x+i\epsilon),$
where we have used the relation $(h_{j})_{|\mathbb{C}^{-}}=\frac{1}{SR((h_{j})_{|\mathbb{C}^{+}})}$ in the first equality and that $\lim_{\epsilon\downarrow 0}O_{j}(x+i\epsilon)\in\mathbb{T}$ in the second. Therefore, it follows from equation (15) that the set $P_{h_{1}}\cup P_{h_{2}}\subset\mathbb{R}$ has full measure.
Now, since $g\in\mathcal{B}(h_{1})\cap\mathcal{B}(h_{2})$ we conclude using Proposition 2.32 that
$\lim_{\epsilon\downarrow 0}g(x+i\epsilon)=\lim_{\epsilon\downarrow 0}g(x-i\epsilon)\text{ for $\lambda$-a.e. $x\in P_{h_{1}}\cup P_{h_{2}}$ or for $\lambda$-a.e. $x\in\mathbb{R}$}.$
This shows that functions $g_{|\mathbb{C}^{+}}$ and $g_{|\mathbb{C}^{-}}$ are pseudocontinuations of each other.
2. The function $g_{|\mathbb{C}^{+}}$ is an element of $\mathcal{H}^{2}(\mathbb{C}^{+})$
It is sufficient to prove that either $O_{1}$ and $O_{2}$ is extreme (Proposition 2.29). Since $O_{2}$ is non-constant, there exists a set $E\subset\mathbb{R}$ such that
$\lambda(E)>0\quad\text{ and }\quad\log(|O_{2}(x)|)<0\quad\text{ for $\lambda$-a.e. $x\in E$}.$
Consequently, $\log(|O_{1}(x)|)=0$, or equivalently $|O_{1}(x)|=1$, holds $\lambda$-a.e. on $E$. Thus $O_{1}$ is extreme.
3. The function $g_{|\mathbb{C}^{-}}$ is an element of $\mathcal{N}^{+}(\mathbb{C}^{-})$.
Let $q_{i}=C^{-1}(h_{i})$ be the associated Herglotz-Nevanlinna function. Moreover, recall that $\mathcal{B}(h_{i})$ is isomorphic to $\mathcal{L}(q_{i})$ via the conjugation
$\mathcal{B}(h_{i})=\frac{1+h_{i}}{\sqrt{2}}\cdot\mathcal{L}(q).$
We have seen in Proposition 2.24 that $\mathcal{L}(q_{i})\subset\mathcal{N}^{+}(\mathbb{C}\setminus\mathbb{R})$. Moreover, it also holds that $(h_{i})_{|\mathbb{C}^{-}}\in\mathcal{N}^{+}(\mathbb{C}^{-})$, since
$(h_{i})_{|\mathbb{C}^{-}}=\frac{1}{SR(O_{i})}$
is an outer function. This means that $g_{|\mathbb{C}^{-}}$ is an element of $\mathcal{N}^{+}(\mathbb{C}^{-})$ because $\mathcal{N}^{+}(\mathbb{C}^{-})$ is closed under multiplications.
Combining the three statements above with Proposition 2.11 yields that $g=0$. ∎
The final argument missing for the proof of Theorem 3.3 is the following rather basic observation:
Lemma 4.4.
Let $g\in\mathcal{N}(\mathbb{C}^{+})\setminus\mathcal{N}^{+}(\mathbb{C}^{+})$. Furthermore, assume that there are two inner functions $W_{1}$ and $W_{2}$ such that $g\cdot W_{1}$ and $g\cdot W_{2}$ are elements of $\mathcal{N}^{+}(\mathbb{C}^{+})$. Then there exists a non-trivial inner function which divides both $W_{1}$ and $W_{2}$.
Proof.
We can write the function $g$ as
$g=\frac{V_{1}}{V_{2}}\cdot F$
where $V_{1}$ and $V_{2}$ are inner functions such that $V_{2}$ is non trivial and $V_{1}$ and $V_{2}$ are relatively prime, and $F$ is an outer function. In order for $g\cdot W_{1}$ or $g\cdot W_{2}$ to be an element of $\mathcal{N}^{+}(\mathbb{C}^{+})$, the inner function $W_{1}$ or $W_{2}$ has to cancel out $V_{2}$. This means that $V_{2}$ divides both $W_{1}$ and $W_{2}$. ∎
Proof of Theorem 3.3.
First assume that there exists a non-trivial function $m\in\mathcal{S}_{0}$ that divides both $h_{1}$ and $h_{2}$. Let $m_{1},m_{2}\in\mathcal{S}_{0}$ be such that $m\cdot m_{i}=h_{i}$. It is straightforward to verify that
$s_{h_{i}}(\zeta,w)=s_{m}(\zeta,w)+m(\zeta)\cdot s_{m_{i}}(\zeta,w)\cdot\overline{m(w)}.$
Consequently, we see that the kernel $s_{h_{i}}$ is the sum of two positive definite kernels. In this case, it is well known [5, Part 1, Section 6], that
$\mathcal{B}(h_{i})=\mathcal{B}(m)+m\cdot\mathcal{B}(m_{i}),$
where the sum is not necessarily direct. Thus, $\mathcal{B}(m)\subset\mathcal{B}(h_{1})\cap\mathcal{B}(h_{2})$ holds, which concludes the proof of this direction.
Conversely, let $h_{1}$ and $h_{2}$ be elements of $\mathcal{S}_{0}$ such that $h_{1}$ and $h_{2}$ are relatively prime. We can decompose $(h_{i})_{|\mathbb{C}^{+}}$ as $(h_{i})_{|\mathbb{C}^{+}}=V_{i}\cdot O_{i}$, where $V_{i}$ is an inner and $O_{i}$ is an outer function. The same reasoning as above yields
$\mathcal{B}(h_{i})=\mathcal{B}(o_{i})+o_{i}\cdot\mathcal{B}(v_{i}),$
where $o_{i}$ and $v_{i}$ denote the functions in $\mathcal{S}_{0}$ generated by $O_{i}$ and $V_{i}$ respectively.
Let $g\in\mathcal{B}(h_{1})\cap\mathcal{B}(h_{2})$. Then we can write $g$ as a sum
$g=f_{1}+o_{1}\cdot k_{1}=f_{2}+o_{2}\cdot k_{2},\quad\text{ with}f_{i}\in\mathcal{B}(o_{i}),\ k_{i}\in\mathcal{B}(v_{i}).$
We have seen in the proof of Lemma 4.3 that $(f_{i})_{|\mathbb{C}^{-}}\in\mathcal{N}^{+}(\mathbb{C}^{-})$ and that $(o_{i})_{|\mathbb{C}^{-}}\in\mathcal{N}^{+}(\mathbb{C}^{-})$. From Proposition 4.2 we know that $(k_{i})_{|\mathbb{C}^{-}}$ is either trivial or $(k_{i})_{|\mathbb{C}^{-}}\notin\mathcal{N}^{+}(\mathbb{C}^{-})$. In the latter case, it then holds that
$(k_{i})_{|\mathbb{C}^{-}}\cdot SR(V_{i})\in\mathcal{N}^{+}(\mathbb{C}^{-}),\ i\in\{1,2\}.$
First assume that $g_{|\mathbb{C}^{-}}\in\mathcal{N}^{+}(\mathbb{C}^{-})$. Then it follows that both $(k_{1})_{|\mathbb{C}^{-}}$ and $(k_{2})_{|\mathbb{C}^{-}}$ have to be trivial, which in turn implies that $g\in\mathcal{B}(o_{1})\cap\mathcal{B}(o_{2})$. However, their intersection is trivial by Lemma 4.3.
Now assume that $g_{|\mathbb{C}^{-}}\notin\mathcal{N}^{+}(\mathbb{C}^{-})$. In this case, we observe that both $g_{|\mathbb{C}^{-}}\cdot SR(V_{1})$ and $g_{|\mathbb{C}^{-}}\cdot SR(V_{2})$ are elements of $\mathcal{N}^{+}(\mathbb{C}^{-})$. Consequently, it follows that the inner functions $SR(V_{1})$ and $SR(V_{2})$ and in turn also $V_{1}$ and $V_{2}$ have a nontrivial common divisor by Lemma 4.4. This contradicts our assumption that $h_{1}$ and $h_{2}$ are relatively prime. ∎
4.2 The behaviour on the boundary
Now we give a proof Lemma 3.4, which is restated for convenience here:
Proposition 4.5.
Let $h_{1},h_{2}\in\mathcal{S}_{0}$ such that they are relatively prime. Then the function
$f=\frac{h_{1}}{h_{2}}$
can be analytically continued through an open interval $(a,b)\subset\mathbb{R}$ if and only if both $h_{1}$ and $h_{2}$ can be analytically continued through $(a,b)$.
Proof.
First, assume that both $h_{1}$ and $h_{2}$ can be analytically continued through $(a,b)$. Then $h_{1}$ and $h_{2}$ have absolute value 1 on $(a,b)$, and are in particular not zero. Consequently, the function $f$ can be analytically continued through $(a,b)$ as well.
Conversely, let $f$ be analytically continuable through $(a,b)$. Then it follows from the definition of $h_{1}$ and $h_{2}$ that
$\overline{f(\zeta)}=\frac{1}{f(\overline{\zeta})}\quad\forall\zeta\in\mathbb{C}\setminus\mathbb{R},$
and we conclude that $|f(w)|=1$ on $(a,b)$. Since $h_{1}$ and $h_{2}$ are relatively prime, it follows that $h_{1}$ and $h_{2}$ can be written as
$(h_{1})_{|\mathbb{C}^{+}}(\zeta)$ $=B_{1}(\zeta)\cdot\exp\left(\frac{-i}{\pi}\int_{-\infty}^{\infty}\frac{-i}{\pi}\int_{\mathbb{R}}\left(\frac{1}{t-\zeta}-\frac{t}{1+t^{2}}\right)\mathrm{d}\mu_{1}(t)\right)\cdot\exp(i\alpha_{1}\zeta)$
$(h_{2})_{|\mathbb{C}^{+}}(\zeta)$ $=B_{1}(\zeta)\cdot\exp\left(\frac{-i}{\pi}\int_{-\infty}^{\infty}\frac{-i}{\pi}\int_{\mathbb{R}}\left(\frac{1}{t-\zeta}-\frac{t}{1+t^{2}}\right)\mathrm{d}\mu_{2}(t)\right)\cdot\exp(i\alpha_{2}\zeta)$
where $B_{1}$ and $B_{2}$ are Blaschke products without common zeros, $d\mu_{1}$ and $d\mu_{2}$ are negative measures which are mutually singular, and at least one of $\alpha_{1}$ and $\alpha_{2}$ is zero. Since a function of the form $\exp(i\alpha\zeta)$ is analytic on $\mathbb{C}$ anyway, we can assume without loss of generality that $\alpha_{1}=\alpha_{2}=0$.
Assume, for sake of contradiction, that there exists an element $\zeta\in(a,b)$ such that $\zeta$ is an accumulation point for the zeros of either $B_{1}$ or $B_{2}$. Then, since $B_{1}$ and $B_{2}$ have no common zeros, there exists a sequence of zeros or poles of $f$ in $\mathbb{C}^{+}$ converging to $\zeta$. This contradicts that $f$ is analytic on $(a,b)$. Therefore, $\zeta$ cannot be such an accumulation point, and we conclude that the functions in $\mathcal{S}_{0}$ generated by $B_{1}$ and $B_{2}$ are analytic on $(a,b)$ and do not vanish there (Proposition 2.12).
Therefore, we can assume without loss of generality that $B_{1}=B_{2}=1$, since $f$ is analytic on $(a,b)$ if and only if $f\cdot\frac{B_{2}}{B_{1}}$ is analytic on $(a,b)$. In other words, we assume that the function $f$ is analytic on $\mathbb{C}^{+}\cup(a,b)$ and does not vanish there. This allows us to consider
$q(\zeta)=-i\cdot\log(f)(\zeta)=\frac{1}{\pi}\int_{\mathbb{R}}\left(\frac{1}{t-\zeta}-\frac{t}{1+t^{2}}\right)\mathrm{d}(\mu_{1}-\mu_{2})(t).$
Note that $q$ is also analytic on $\mathbb{C}^{+}\cup(a,b)$. It is known [39, Corollary 4.3], that the measure $\mu_{1}-\mu_{2}$ is given by the weak-(*)(*)( * ) limit of the densities
$p_{y}(t)=\mathrm{Im}(q)(t+iy)=\log{|f(t+iy)|},\quad y>0.$
Since $|f(w)|=1$ on $(a,b)$ and $f$ is analytic on $(a,b)$, it follows that the densities $p_{y}$ go to zero locally uniformly on $(a,b)$. Therefore, we conclude that $(\mu_{1}-\mu_{2})\cdot\mathbbm{1}_{(a,b)}$ is the trivial measure. Since $\mu_{1}$ and $\mu_{2}$ are mutually singular by construction, it follows that also $\mu_{1}\cdot\mathbbm{1}_{(a,b)}$ and $\mu_{2}\cdot\mathbbm{1}_{(a,b)}$ are trivial [23, Section 2]. This means that $h_{1}$ and $h_{2}$ are analytic on $(a,b)$ [45, Theorem 3.20] . ∎
5 Model relations on Krein spaces
The construction of minimal realizations presented in Section 3 defines an interesting class of self-adjoint relations on Krein spaces, which we will call model relations. In this section we describe their structure in detail. Our first Proposition is a simple transformation result, which will turn out useful at a later stage.
Proposition 5.1.
Let $f\in\mathcal{N}_{sym}$ and $(\mathcal{L}(f),A_{f},\phi_{f})$ be the realization constructed in Section 3, where
$\mathcal{L}(f)=\frac{\sqrt{2}}{h_{1}+h_{2}}\cdot(\mathcal{B}(h_{1})\oplus(-\mathcal{B}(h_{2}))$
for a pair of functions $h_{1},h_{2}\in\mathcal{S}_{0}$ which are relatively prime. Under this conjugation $A_{f}$ transforms to
$\tilde{A}_{f}=\left\{(h,g)\in(\mathcal{B}(h_{1})\oplus-\mathcal{B}(h_{2}))^{2}\ |\ \exists c\in\mathbb{C}\vcentcolon g(\zeta)-\zeta\cdot h(\zeta)\equiv c\cdot\left(h_{1}(\zeta)+h_{2}(\zeta)\right)\right\}.$
The advantage of considering $A_{f}$ on $\mathcal{B}(h_{1})\oplus-\mathcal{B}(h_{2})$ is that in this case we are given an explicit fundamental decomposition consisting of two well-studied function spaces.
Proof.
Recall that $A_{f}$ on $\mathcal{L}(q)$ is given by
$A_{f}=\{(h,g)\in\mathcal{L}(q)\times\mathcal{L}(q)\vcentcolon\exists c\in\mathbb{C}\vcentcolon g(\zeta)-\zeta\cdot h(\zeta)\equiv c\}.$
It is then straightforward to verify that $\tilde{A}_{f}$ has the claimed form.
The theory developed in [10] provides a suitable framework to describe the structure of model relations. Following their approach, we introduce partially fundamentally reducible relations:
Definition 5.2.
Let $A$ be a self-adjoint relation on a Krein space $\mathcal{K}$ satisfying $\varrho(A)\neq\emptyset$. Then
  • the relation $A$ is called fundamentally reducible if there exists a fundamental decomposition $\mathcal{K}=\mathcal{K}_{+}[+]\mathcal{K}_{-}$ such that $A$ splits into an orthogonal sum $A=A_{1}\oplus A_{2}$ with respect to this decomposition.
  • the relation $A$ is called partially fundamentally reducible if there exists a fundamental decomposition $\mathcal{K}=\mathcal{K}_{+}[+]\mathcal{K}_{-}$ such that the relations
    $S_{+}\vcentcolon=A\cap(\mathcal{K}_{+}\times\mathcal{K}_{+})\quad\text{ and }\quad S_{-}\vcentcolon=A\cap(\mathcal{K}_{-}\times\mathcal{K}_{-})$
    are closed, simple symmetric operators on the Hilbert spaces $\mathcal{K}_{+}$ and $-\mathcal{K}_{-}$ with deficiency index $(1,1)$.
The structure of fundamentally reducible relations is relatively easy to analyse, since such a relation is also self-adjoint with respect to the induced Hilbert space inner product. As the name suggests, partially fundamentally reducible relations are relatively close to fundamentally reducible ones. This will be specified at a later stage. First, we prove that model relations are indeed of that form:
Theorem 5.3.
Let $A_{f}$ be a model relation corresponding to a function $f$ such that neither $f$ nor $-f$ is a Herglotz-Nevanlinna function. Then $A_{f}$ is partially fundamentally reducible.
Proof.
We assume without loss of generality that $A_{f}$ is given by
$A_{f}=\left\{(h,g)\in(\mathcal{B}(h_{1})\oplus-\mathcal{B}(h_{2}))^{2}|\ \exists c\in\mathbb{C}\vcentcolon g(\zeta)-\zeta\cdot h(\zeta)\equiv c\cdot\left(h_{1}(\zeta)+h_{2}(\zeta)\right)\right\}.$
Consider an element $(h,g)\in A_{f}\cap(\mathcal{B}(h_{1})\times\mathcal{B}(h_{1}))$. Then there exists a constant $c\in\mathbb{C}$ such that
$g(\zeta)-\zeta\cdot h(\zeta)\equiv c\cdot\left(h_{1}(\zeta)+h_{2}(\zeta)\right).$
Applying $D_{w}$ to the left-hand side, for a suitable $w$, and using its multiplication identity, we obtain
$D_{w}(g-\zeta\cdot f)=D_{w}(g)-D_{w}(\zeta\cdot f)=D_{w}(g)-w\cdot D_{w}(f)-f\cdot D_{w}(\zeta)=D_{w}(g)-w\cdot D_{w}(f)-f\cdot 1.$
This is an element of $\mathcal{B}(h_{1})$, because both $f$ and $g$ are and $D_{w}$ acts boundedly on $\mathcal{B}(h_{1})$. Consequently, it follows that also
$D_{w}(c\cdot\left(h_{1}+h_{2}\right))=c\cdot(D_{w}(h_{1})+D_{w}(h_{2}))\in\mathcal{B}(h_{1}).$
However, we know that $D_{w}(h_{i})\in\mathcal{B}(h_{i})$. This means that $D_{w}(c\cdot\left(h_{1}+h_{2}\right))$ is either the zero function if $c=0$, or not an element of $\mathcal{B}(h_{1})$. Consequently, we arrive at
$A_{f}\cap(\mathcal{B}(h_{1})\times\mathcal{B}(h_{1}))=\left\{(h,g)\in\mathcal{B}(h_{1})\times\mathcal{B}(h_{1})\ |g(\zeta)-\zeta\cdot h(\zeta)\equiv 0\right\}.$
This means that $A_{f}\cap(\mathcal{B}(h_{1})\times\mathcal{B}(h_{1}))$ is the multiplication operator with the independent variable on $\mathcal{B}(h_{1})$, which is a simple, symmetric operator with deficiency index $(1,1)$. The second relation $A_{f}\cap(\mathcal{B}(h_{2})\times\mathcal{B}(h_{2}))$ is treated analogously. ∎
The next theorem is, in principle, a special case of [10, Theorem 4.7]. The primary distinction in our context is that the symmetric operators we consider are not necessarily densely defined, an assumption that was not crucial to their argument. In our specific situation, this result can be proven in a more elementary manner using the multiplication identity of $D_{w}$.
Theorem 5.4.
Let $f\in\mathcal{N}_{sym}$ such that neither $f$ nor $-f$ is a Herglotz-Nevanlinna function. Moreover, consider the realization $(\mathcal{L}(f),A_{f},\phi_{f})$ constructed in Section 3, where
$\mathcal{L}(f)=\frac{\sqrt{2}}{h_{1}+h_{2}}\cdot(\mathcal{B}(h_{1})\oplus(-\mathcal{B}(h_{2}))$
for a pair of functions $h_{1},h_{2}\in\mathcal{S}_{0}$ which are relatively prime. Finally, set
$q_{1}(\zeta)=i\cdot\frac{1-h_{1}(\zeta)}{1+h_{1}(\zeta)}\quad k_{1}(\zeta)\vcentcolon=\frac{1+h_{1}(\zeta)}{h_{1}(\zeta)+h_{2}(\zeta)}$
$q_{2}(\zeta)=-i\cdot\frac{1+h_{2}(\zeta)}{1-h_{2}(\zeta)}\quad k_{2}(\zeta)\vcentcolon=\frac{1-h_{2}(\zeta)}{h_{1}(\zeta)+h_{2}(\zeta)}$
$\phi(w)=k_{1}\cdot D_{w}(q_{1})+k_{2}\cdot D_{w}(q_{2}).$
Then $\phi(w)\in\mathcal{L}(f)$ for every $w\in\varrho(A_{f})\cap\mathbb{C}\setminus\mathbb{R}$, and there exists a fundamentally reducible relation $\tilde{A}$ such that
$(A_{f}-w)^{-1}=(\tilde{A}-w)^{-1}-\frac{[\ \cdot\ ,\phi(\overline{w})]}{q_{1}(w)+q_{2}(w)}\cdot\phi(w)\quad\forall w\in\varrho(A)\cap\varrho(\tilde{A}).$ (16)
Proof.
First, we note that the functions $q_{1}$ and $q_{2}$ are related to $h_{1}$ and $h_{2}$ via the Cayley transform. Specifically, we have:
$q_{1}=C(h_{1})\text{ and }\frac{1}{q_{2}}=C(h_{2}).$
The functions $C(h_{i}),\ i\in\{1,2\}$ are Herglotz-Nevanlinna functions. This means that both $q_{1}$ and $-q_{2}$ are Herglotz-Nevanlinna functions as well. Moreover, we can represent $\mathcal{L}(f)$ as
$\mathcal{L}(f)=\frac{\sqrt{2}}{h_{1}+h_{2}}\cdot(\mathcal{B}(h_{1})\oplus(-\mathcal{B}(h_{2}))=k_{1}\cdot\mathcal{L}(q_{1})\oplus-(k_{2}\cdot\mathcal{L}(-q_{2})).$
This representation is a consequence of Theorem 2.27, according to which it holds that
$\begin{split}\frac{\sqrt{2}}{h_{1}+h_{2}}\cdot\mathcal{B}(h_{1})&=\frac{1+h_{1}}{h_{1}+h_{2}}\cdot\mathcal{L}(q_{1})=k_{1}\cdot\mathcal{L}(q_{1})\\ \frac{\sqrt{2}}{h_{1}+h_{2}}\cdot\mathcal{B}(h_{2})&=\frac{1+h_{2}}{h_{1}+h_{2}}\cdot\mathcal{L}\left(\frac{1}{q_{2}}\right)=\frac{1+h_{2}}{h_{1}+h_{2}}\cdot\frac{1}{q_{2}}\cdot\mathcal{L}(-q_{2})=k_{2}\cdot\mathcal{L}(-q_{2}).\end{split}$ (17)
Here, we have used the identity $\mathcal{L}\left(\frac{1}{q_{2}}\right)=\frac{1}{q_{2}}\cdot\mathcal{L}(-q_{2})$, which was established in Section 3.1. Additionally, we utilized that unimodular factors do not affect conjugations.
Now consider the self-adjoint relations $A_{1}$ and $A_{2}$ on $\mathcal{L}(q_{1})$ and $\mathcal{L}(-q_{2})$ given by $(A_{i}-w)^{-1}=D_{w}$. We define the desired fundamentally reducible self-adjoint relation $\tilde{A}\subset\mathcal{L}(f)\times\mathcal{L}(f)$ as
$\tilde{A}=\left(k_{1}\cdot A_{1}\cdot\frac{1}{k_{1}}\right)\oplus\left(k_{2}\cdot A_{1}\cdot\frac{1}{k_{2}}\right),$
where the “multiplications” denote the isomorphisms induced by the conjugations detailed in (17). Next, let $k_{1}\cdot g$ belong to $k_{1}\cdot\mathcal{L}(q_{1})$. We compute
$((A_{f}-w)^{-1}-(\tilde{A}-w)^{-1})(k_{1}\cdot g)$ $=D_{w}(k_{1}\cdot g)-\left(k_{1}\cdot D_{w}\cdot\frac{1}{k_{1}}\right)(k_{1}\cdot g)$
$=k_{1}\cdot D_{w}(g)+g(w)\cdot D_{w}(k_{1})-k_{1}\cdot D_{w}(g)=g(w)\cdot D_{w}(k_{1}).$
We also observe that
$[k_{1}\cdot g,\phi(\overline{w})]_{\mathcal{L}(f)}=[k_{1}\cdot g,k_{1}\cdot N_{q_{1}}(\cdot,w)]_{\mathcal{L}(f)}=[g,N_{q_{1}}(\cdot,w)]_{\mathcal{L}(q_{1})}=q(w),$
where we have used the conjugation given by (17). Therefore, to establish (16), we only need to demonstrate that
$D_{w}(k_{1})\cdot(q_{1}(w)+q_{2}(w))+\phi(w)=0.$
To this end, we first note that $k_{1}=1+k_{2}$, which means that $D_{w}(k_{1})=D_{w}(k_{2})$. Consequently, it follows that
$D_{w}(k_{1})\cdot(q_{1}(w)+q_{2}(w))+\phi(w)$ $=D_{w}(k_{1})\cdot q_{1}(w)+k_{1}\cdot D_{w}(q_{1})+D_{w}(k_{2})\cdot q_{2}(w)+k_{2}\cdot D_{w}(q_{2})$
$=D_{w}(k_{1}\cdot q_{1})+D_{w}(k_{2}\cdot q_{2})=D_{w}(k_{1}\cdot q_{1}+k_{2}\cdot q_{2})=0,$
where we have used the multiplication identity of $D_{w}$ and
$k_{1}\cdot q_{1}+k_{2}\cdot q_{2}=-i.$
A similar calculation shows that (16) is also valid for functions in the negative part of $\mathcal{L}(f)$, which is $k_{2}\cdot\mathcal{L}(-q_{2})$. This concludes our proof. ∎
A similar idea also allows us to prove the “converse“ of Theorem 5.3. This, and the preceding results, are summarized in the following theorem:
Theorem 5.5.
Let $q\in\mathcal{N}_{sym}$. Then $q$ has a minimal realization $(\mathcal{K},A,v)$ where $A$ is partially fundamentally reducible.
Conversely, let $A$ be a partially fundamentally reducible relation on $\mathcal{K}$. Then $\sigma(A)\cap(\mathbb{C}\setminus\mathbb{R})$ is a discrete set in $\mathbb{C}\setminus\mathbb{R}$. Moreover, if $(\mathcal{K},A,v)$ is a realization of a function $q$ with base point $\zeta_{0}\in\varrho(A)$, then $q$ is an element of $\mathcal{N}_{sym}$.
Proof.
Let $A$ be a partially fundamentally reducible relation on $\mathcal{K}$. Then there exists a fundamental decomposition $\mathcal{K}=\mathcal{K}_{+}[+]\mathcal{K}_{-}$ such that the relations
$S_{+}\vcentcolon=A\cap(\mathcal{K}_{+}\times\mathcal{K}_{+})\quad\text{ and }\quad S_{-}\vcentcolon=A\cap(\mathcal{K}_{-}\times\mathcal{K}_{-})$
are closed, simple symmetric operators on the Hilbert spaces $\mathcal{K}_{+}$ and $-\mathcal{K}_{-}$ with deficiency index $(1,1)$.
Let $(\Gamma_{0}^{+},\Gamma_{1}^{+},\mathbb{C})$ and $(\Gamma_{0}^{-},\Gamma_{1}^{-},\mathbb{C})$ be arbitrary but fixed boundary triples for $S_{+}$ and $S_{-}$, which exist by Proposition 2.37. Moreover, let $q_{+}$ and $q_{-}$ denote the Weyl-functions, and $y_{+}$ and $y_{-}$ denote the Weyl-solutions corresponding to these boundary triples. Note that $q_{+}$ and $q_{-}$ are Herglotz-Nevanlinna functions. We also define self-adjoint relations $A_{+}$ and $A_{-}$ by
$A_{+}=\mathrm{ker}(\Gamma_{0}^{+})\quad\text{ and }A_{-}=\mathrm{ker}(\Gamma_{0}^{-}).$
Now consider the closed and symmetric operator $S_{+}\oplus S_{-}$ on the Krein space $\mathcal{K}$. It is clear by construction that both $A$ and $A_{+}\oplus A_{-}$ are self-adjoint extension of $S_{+}\oplus S_{-}$. Moreover, the triple
$(\Gamma_{0}^{+}\oplus(-\Gamma_{0}^{-}),\Gamma_{1}^{+}\oplus\Gamma_{1}^{-},\mathbb{C}^{2})$
is the boundary triple for $S_{+}\oplus S_{-}$ satisfying $\mathrm{ker}((\Gamma_{0}^{+}\oplus(-\Gamma_{0}^{-}))=A_{0}\oplus A_{1}$. The respective Weyl solution $\gamma$ is given by $\gamma=\gamma_{+}\oplus\gamma_{-}$, and the respective Weyl function by
$m(\zeta)=\begin{pmatrix}q_{+}(\zeta)&0\\ 0&-q_{-}(\zeta)\end{pmatrix}.$
This allows us to use Krein‘s resolvent formalism [14, Corollary 2.4]. More precisely, there exists two $2\times 2$ matrices $G=(g_{ij})$ and $H=(h_{ij})$ such that
$\zeta\in\varrho(A)\Leftrightarrow 0\in\varrho(G-m(\zeta)\cdot H)\Leftrightarrow\mathrm{det}(G-m(\zeta)\cdot H)\neq 0.$
We can explicitly compute the determinant as
$\mathrm{det}(G-m(\zeta)\cdot H)$ $=\mathrm{det}\left(\begin{pmatrix}g_{11}-q_{+}(\zeta)\cdot h_{11}&g_{12}-q_{+}(\zeta)\cdot h_{12}\\ g_{21}+q_{-}(\zeta)\cdot h_{21}&g_{22}+q_{-}(\zeta)\cdot h_{22}\end{pmatrix}\right)$
$=(g_{11}-q_{+}(\zeta)\cdot h_{11})\cdot(g_{22}+q_{-}(\zeta)\cdot h_{22})-(g_{21}+q_{-}(\zeta)\cdot h_{21})\cdot(g_{12}-q_{+}(\zeta)\cdot h_{12}).$
This is an analytic function in $\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$, which means that $\sigma(A)\cap(\mathbb{C}\setminus\mathbb{R})$ is a discrete set in $\mathbb{C}\setminus\mathbb{R}$. Finally, Krein’s resolvent formula takes the form
$(A-\zeta)^{-1}=(A_{0}-\zeta)^{-1}+\gamma(\zeta)(H(G-m(\zeta)\cdot H)^{-1})\gamma^{+}(\overline{\zeta})\quad\zeta\in\varrho(A),$
which allows us to calculate
$\begin{split}\left[\left(I+(\zeta-\zeta_{0})(A-\zeta)^{-1}\right)v,v\right]_{\mathcal{K}}&=[(I+(\zeta-\zeta_{0})(A_{0}-\zeta)^{-1}v,v]_{\mathcal{K}}\\ &+[\gamma(\zeta)(H(G-m(\zeta)\cdot H)^{-1})\gamma^{+}(\overline{\zeta})v,v]_{\mathcal{K}}.\end{split}$ (18)
In what follows, our strategy is to show that all occurring functions are of bounded type. Let $v=v_{+}+v_{-}$ be the decomposition of $v$ with respect to the fundamental decomposition $\mathcal{K}_{+}[+]\mathcal{K}_{-}$. Then the first summand simplifies to
$[(I+(\zeta-\zeta_{0})(A_{0}-\zeta)^{-1}v,v]_{\mathcal{K}}$ $=[(I+(\zeta-\zeta_{0})(A_{+}-\zeta)^{-1}v_{+},v_{+}]_{\mathcal{K}_{+}}$
$+[(I+(\zeta-\zeta_{0})(A_{-}-\zeta)^{-1}v_{-},v_{-}]_{\mathcal{K}_{-}}=s_{1}-s_{2},$
where $s_{1}$ and $s_{2}$ are Herglotz-Nevanlinna functions and therefore elements of $\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$. Here, we have used that $A_{0}$ decomposes into an orthogonal sum $A_{0}=A_{+}\oplus A_{-}$ with respect to the fundamental decomposition $\mathcal{K}=\mathcal{K}_{+}[+]\mathcal{K}_{-}$.
Now we turn our attention to the second summand in (18). It was shown in Proposition 2.37 that
$\zeta\mapsto(\gamma_{+}(\overline{\zeta}))^{+}(v_{+})$ $=[v_{+},\gamma_{+}(\overline{\zeta})]_{\mathcal{K}_{+}}\in\mathcal{L}(q_{1})\subset\mathcal{N}(\mathbb{C}\setminus\mathbb{R}),$
$\zeta\mapsto(\gamma_{-}(\overline{\zeta}))^{+}(v_{-})$ $=[v_{-},\gamma_{-}(\overline{\zeta})]_{\mathcal{K}_{-}}\in\mathcal{L}(-q_{2})\subset\mathcal{N}(\mathbb{C}\setminus\mathbb{R}).$
Moreover, we compute $H(G-m(\zeta)\cdot H)^{-1}$ using the formula for the inverse of a $2\times 2$ matrix as
$H(G-m(\zeta)\cdot H)^{-1}$ $=\mathrm{det}(G-m(\zeta)\cdot H)\cdot H\cdot\begin{pmatrix}g_{22}+q_{-}(\zeta)\cdot h_{22}&-g_{21}-q_{-}(\zeta)\cdot h_{21}\\ -g_{12}+q_{+}(\zeta)\cdot h_{12}&g_{11}-q_{+}(\zeta)\cdot h_{11}\end{pmatrix}$
:$\vcentcolon=\begin{pmatrix}l_{11}(\zeta)&l_{12}(\zeta)\\ l_{21}(\zeta)&l_{22}(\zeta)\end{pmatrix}.$
Since $\mathrm{det}(G-m(\zeta)\cdot H)$ is an element of $\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$ and the entries of $H$ are constants, it follows that the entries $(l_{ij}(\zeta))$ are also elements of $\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$. We conclude that
$\gamma(\zeta)(H(G-m(\zeta)\cdot H)^{-1})\gamma^{+}(\overline{\zeta})v$ $=\gamma(\zeta)\begin{pmatrix}l_{11}(\zeta)&l_{12}(\zeta)\\ l_{21}(\zeta)&l_{22}(\zeta)\end{pmatrix}\cdot\begin{pmatrix}[v_{+},\gamma_{+}(\overline{\zeta})]\\ [v_{-},\gamma_{-}(\overline{\zeta})]\end{pmatrix}$
$=(l_{11}(\zeta)\cdot[v_{+},\gamma_{+}(\overline{\zeta})]+l_{12}(\zeta)\cdot[v_{-},\gamma_{-}(\overline{\zeta})])\cdot\gamma_{+}(\zeta)$
$+(l_{21}(\zeta)\cdot[v_{+},\gamma_{+}(\overline{\zeta})]+l_{22}(\zeta)\cdot[v_{-},\gamma_{-}(\overline{\zeta})])\cdot\gamma_{-}(\zeta)$
$=s_{3}(\zeta)\cdot\gamma_{+}(\zeta)+s_{4}(\zeta)\cdot\gamma_{-}(\zeta),$
where $s_{3}$ and $s_{4}$ are functions of the class $\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$. Consequently, it follows that
$[\gamma(\zeta)(H(G-m(\zeta)\cdot H)^{-1})\gamma^{+}(\zeta)v,v]_{\mathcal{K}}$ $=[s_{3}(\zeta)\cdot\gamma_{+}(\zeta)+s_{4}(\zeta)\cdot\gamma_{-}(\zeta),v]_{\mathcal{K}}$
$=s_{3}(\zeta)\cdot\overline{[v_{+},\gamma_{+}(\zeta)]}_{\mathcal{K}_{+}}+s_{4}(\zeta)\cdot\overline{[v_{-},\gamma_{-}(\zeta)]}_{\mathcal{K}_{-}}\in\mathcal{N}(\mathbb{C}\setminus\mathbb{R}),$
where we have used that $\overline{[v_{+},\gamma_{+}(\zeta)]}_{\mathcal{K}_{+}}$ and $\overline{[v_{-},\gamma_{-}(\zeta)]}_{\mathcal{K}_{-}}$ belong to $\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$.
Originally, we had assumed that $(\mathcal{K},A,v)$ is a realization of a function $q$, i.e. that
$q(\zeta)=\overline{q(\zeta_{0})}+(\zeta-\overline{\zeta_{0}})\cdot\left[\left(I+(\zeta-\zeta_{0})(A-\zeta)^{-1}\right)v,v\right]_{\mathcal{K}}\quad\forall\zeta\in\varrho(A).$
Our preceding analysis has shown that the inner product coincides with a function in $\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$ on $\varrho(A)$, which means that $q\in\mathcal{N}(\mathbb{C}\setminus\mathbb{R})$. This concludes the proof. ∎
6 Examples
6.1 Generalized Nevanlinna functions
There are several conceptually different ways to construct minimal realizations for generalized Nevanlinna functions. For example, the construction in [15] is based on reproducing kernel Pontryagin spaces, while the one in [20] relies on a certain factorization result. In this article, we have used the classical theory of model spaces and related function spaces. This function-theoretic approach allows us to give a short reproof of the following result [46, Theorem 1.1]:
Theorem 6.1.
Let $f$ be a Generalized Nevanlinna function with index $\kappa$. For any given $w\in\mathbb{C}^{-}$, the equation $f(\zeta)-w=0$ has exactly $\kappa$ solutions on $\mathbb{C}^{+}$ (counted with multiplicities).
Proof.
Let $f$ be represented as
$f=i\cdot\frac{h_{2}-h_{1}}{h_{2}+h_{1}},$
where $h_{1}\in\mathcal{S}_{0}$ and $h_{2}\in\mathcal{S}_{0}$ are as in Proposition 3.2. The index $\kappa$ coincides with the dimension of the negative part of $\mathcal{L}(f)$, which means that $\kappa=\mathrm{dim}(\mathcal{B}(h_{2}))$. Consequently, the function $h_{2}\in\mathcal{S}_{0}$ is generated by a finite Blaschke product with $\kappa$ zeros see Section 2.5. However, the zeros of $h_{2}$ in $\mathbb{C}^{+}$ correspond exactly to the zeros of $f+i=0$ in $\mathbb{C}^{+}$(counted with multiplicities).
In summary, we have proven our claim for $w=-i$. We extend this to arbitrary points in the lower-half plane in the following way: First, note that the function $a\cdot f+b,\ a,b\in\mathbb{R}_{+}$ has the same index as $f$. Therefore, we can apply our result for $w=-i$ to this function, and we infer that the equation
$(a\cdot f+b)+i=0\Leftrightarrow f-\frac{i+b}{-a}=0,\quad a,b\in\mathbb{R}_{+}$
has exactly $\kappa$ solutions on $\mathbb{C}^{+}$ (counted with multiplicities). Finally, it is clear that we can write any $w\in\mathbb{C}^{-}$ as $w=\frac{i+b}{-a}$ with $a,b\in\mathbb{R}_{+}$. ∎
6.2 Real complex functions
Next we discuss real complex function which have attracted substantial interest in the recent past, see [24], [28] and the survey [26]. They are defined as follows:
Definition 6.2.
Let $f\in\mathcal{N}_{sym}$. Then $f$ is a real complex function function if $f_{|\mathbb{C}^{+}}$ has real boundary values $\lambda$-a.e.
We have seen in Section 4 that such a function $f$ can be represented as
$f=i\cdot\frac{h_{2}-h_{1}}{h_{2}+h_{1}},$
where $h_{1}$ and $h_{2}$ are generated by inner functions $V_{1}$ and $V_{2}$. In this case, the associated model spaces $\mathcal{H}(V_{1})$ and $\mathcal{H}(V_{2})$ intersect trivially, since $h_{1}$ and $h_{2}$ are relatively prime and in turn also $V_{1}$ and $V_{2}$ are relatively prime. Consequently, the restriction operator
$\mathrm{Res}\vcentcolon\mathcal{B}(h_{1})\oplus-\mathcal{B}(h_{2})\rightarrow\mathcal{H}(V_{1})\oplus-\mathcal{H}(V_{2}),\quad f\mapsto f_{|\mathbb{C}^{+}}$
is an isometric isomorphism. This means that its realization acts, up to an isomorphism, on a direct outer sum of Model spaces. As already mentioned, the structure of model spaces has been the subject of extensive research and is very well understood.
Here, we want to give one example of how our construction takes the structure of the underlying function $f$ into account. If $f$ is a real complex function, then $f_{|\mathbb{C}^{+}}$ and $f_{|\mathbb{C}^{-}}$ are pseudocontinuations of each other, and this property is inherited by the space $\mathcal{L}(f)$. Indeed, since this is true for any function $g\in\mathcal{B}(h_{1})\oplus-\mathcal{B}(h_{2})$ and for the function
$\frac{1}{h_{1}+h_{2}},$
it is also true for the function
$\frac{1}{h_{1}+h_{2}}\cdot g\in\frac{1}{h_{1}+h_{2}}\cdot(\mathcal{B}(h_{1})\oplus-\mathcal{B}(h_{2}))=\mathcal{L}(f).$
6.3 The extended Nevanlinna class
Originally, we were interested in the following function class, which was introduced in [13]:
Definition 6.3.
Let $f\in\mathcal{N}_{sym}$. Then $f$ is in the extended Nevanlinna class if $f$ is a product of the form $f=g\cdot m$, where $g$ is a Herglotz-Nevanlinna function and $m$ is a density function. Here, a density function is a real complex function with non-negative boundary values.
The extended Nevanlinna class generalizes the notion of generalized Nevanlinna functions in a natural way by allowing the density factor $m$ to be non-rational. On the other hand, this class also generalizes the class of real complex functions. Indeed, it was shown in [13, Theorem 3.1] that any real complex function $f$ is a product of the form
$f=g\cdot m$
where $g$ is a Herglotz-Nevanlinna function with a real boundary values and $m$ is a density function. In any case, we can represent $f$ as
$f=i\cdot\frac{h_{2}-h_{1}}{h_{2}+h_{1}}$
where $h_{2}$ is generated by an inner function.